### v8

# An interacting particle consensus method for constrained global optimization

Abstract.
This paper presents a particle-based optimization method designed for addressing minimization problems with equality constraints, particularly in cases where the loss function exhibits non-differentiability or non-convexity. The proposed method combines components from consensus-based optimization algorithm with a newly introduced forcing term directed at the constraint set. A rigorous mean-field limit of the particle system is derived, and the convergence of the mean-field limit to the constrained minimizer is established. Additionally, we introduce a stable discretized algorithm and conduct various numerical experiments to demonstrate the performance of the proposed method.
Key words and phrases:
constrained optimization; gradient-free methods; nonconvex optimization; consensus-based-optimization; asymptotic convergence analysis; mean-field limit
2020 Mathematics Subject Classification:
90C56; 65C35; 35Q70; 82C22; 35Q84
1 Mathematical Institute, University of Oxford, United Kingdom. (carrillo@maths.ox.ac.uk)
2 School of Mathematical Sciences, Institute of Natural Sciences, MOE-LSC, Shanghai Jiao Tong University, Shanghai, 200240, P. R. China. (shijin-m@sjtu.edu.cn)
3 Department of Mathematics, University of California-San Diego, La Jolla, California 92093, USA. (haz053@ucsd.edu)
4 Department of Statistics and Data Science, University of California, Los Angeles, USA (yuz244@ucsd.edu)
All authors contributed equally to this work and are listed alphabetically.
$\dagger$ Corresponding author.
1. Introduction
In this paper, we are concerned with the following minimization problem with equality constraint,
$\min_{v\in\mathbb{R}^{d}}\quad$ $\mathcal{E}(v)$
s.t. ${g(v)=0.}$
The above optimization problems have widespread application across various domains. For example, in supply chain optimization, equality constraints play a pivotal role in maintaining a balance between demand and supply [37]; astronomers employ constrained optimization to calculate spacecraft trajectories, adhering to the laws of physics and orbital equations [15, 52]; in structural design, engineers optimize dimensions of beams, columns, or trusses while ensuring that the structural equilibrium equations are satisfied as equality constraints [27]. In this paper we deal with the cases when the objective function can be non-convex and non-differentiable.
Traditional algorithms like the Lagrange Multipliers [46] and the Alternating Direction Method of Multipliers (ADMM) [51] lack guarantees of converging to the global constrained minimizer when dealing with non-convex or non-differentiable loss functions $\mathcal{E}(v)$. A new framework is required to effectively handle such cases, and recently, a class of gradient-free methods called consensus-based optimization (CBO) methods [4, 12, 41, 48] have emerged as promising approaches for handling non-convex and non-differentiable loss functions. Motivated by the well-known Laplace’s principle [7, 17, 38], they are decentralized and gradient-free algorithms that leverage the power of information sharing and cooperation among individual particles. However, it is important to highlight that much of the existing work has focused on the unconstrained case, see [12, 13, 16, 23, 24, 25, 29, 30, 31, 32, 44, 43, 49]. We refer the readers to survey articles [14, 28, 50] for a more detailed and complete summary. For the efficient implementation of CBO, we refer to the CBX packages developed for Python and Julia [4].
Limited work has been done for the constrained case. The primary challenge lies in reconciling the CBO model’s tendency to drive agents towards the global minimizer with the need for agents to remain within the constraint set and converge to the constrained minimizer. Currently, there are mainly two approaches. One involves projection onto the hypersurface [2, 19, 20, 21]. However, this method requires computing the derivatives of the distance function $\operatorname{dist}(\Gamma,v)=\inf\{\|v-u\|_{2}\mid u\in\Gamma\}$ with $\Gamma$ representing the constraint set. Extending this method to handle general multiple equality constraints is not straightforward. In cases where the constraint set $\Gamma$ is complicated, this computation about $\operatorname{dist}(\Gamma,v)$ becomes infeasible. Another method introduces constraints as a penalization term in the objective function [8, 14], transforming it into an unconstrained problem for CBO. However, the convergence is sensitive to the landscape of the objective function and the penalization constant, which makes it difficult to achieve high accuracy.
In this paper, we introduce a third strategy for constrained CBO along with convergence analysis and numerical experiments. Instead of performing projection onto the constraint set or adding penalty terms, we propose a novel approach that combines the classical unconstrained CBO algorithm with gradient descent on the function $G(v)={g^{2}(v)}$, serving as a forcing term to the constraint set. Importantly, we do not require the differentiability of the target function $\mathcal{E}$ and only need a mild differentiability condition on $G$. Compared with the other two constrained CBO methods (projected CBO and penalized CBO), our method applies to general equality constraints, achieves faster convergence, and has consistently more stable performance as shown in Figures 1 and 2.
1.1. Contributions
Our main contributions are threefold. Firstly, we introduce a new CBO-based method for solving constrained optimization problems, with possibly non-convex and non-differentiable objective functions. This method can accommodate a wide range of equality constraints, including the ability to handle multiple constraints concurrently. Secondly, we provide rigorous theoretical guarantees for the continuous-in-time model of the proposed method. Specifically, we establish the mean-field limit of the method and provide a detailed analysis of its convergence behavior within this limit, using a new quantitative Laplace Principle that differs from the unconstrained version in [23]. This principle enables us to quantify the contributions of the consensus dynamics and gradient descent, providing a framework for handling constraint sets. See Remark 5 for the details. Thirdly, we present a stable discretized algorithm designed to approximate the dynamics of the continuous-in-time model efficiently. Notably, this algorithm handles the stiff term of order $O(\epsilon^{-1})$ without requiring the time step to approach zero when $\epsilon$ becomes very small.
1.2. Organizations
The paper is structured as follows. Section 2 provides an introduction to the continuous-in-time stochastic differential equations, which serves as the model for the proposed method. Following that, Section 3 studies the well-posedness of the introduced SDEs and explores their mean-field limit. In Section 4, we analyze the convergence properties of the method by establishing the long-time behavior of the mean-field limit. This includes demonstrating, under appropriate assumptions, the convergence of the mean-field limit model to the constrained minimizer. Section 5 details the implementation of the algorithm, accompanied by a series of numerical experiments showcasing its performance. Finally, Section 6 offers a comprehensive summary of the findings presented in this paper.
1.3. Notations
We use $\mathcal{C}^{k}_{b}(\mathbb{R}^{d})$ and $\mathcal{C}^{k}_{c}(\mathbb{R}^{d})$ to denote the space of $k$-times continuous differentiable functions defined on $\mathbb{R}^{d}$ that are bounded and compactly supported respectively. The space $\mathcal{C}_{*}^{2}$ is defined as
$\mathcal{C}^{2}_{*}(\mathbb{R}^{d}):=\left\{\begin{array}[]{c}\phi\in\mathcal{C}^{2}(\mathbb{R}^{d})|\text{ }|\partial_{x_{k}}\phi(x)|\leq C(1+|x_{k}|)\\ \text{ and }\\ \sup_{x\in\mathbb{R}^{d}}|\partial_{x_{k}x_{k}}\phi(x)|<\infty\text{ for all $k=1,2,...,d$}\end{array}\right\}.$
When $X$ and $Y$ are topological spaces, we use $\mathcal{C}(X,Y)$ to denote the space of continuous functions mapping from $X$ to $Y$. When $X$ is a topological space, $\mathcal{P}(X)$ denotes the space of all the Borel probability measure, which is equipped with the Levy-Prokhorov metric. Given $1\leq p<\infty$, $\mathcal{P}_{p}(\mathbb{R}^{d})$ is the collection of all probability measures on $\mathbb{R}^{d}$ with finite $p$-th moment, which is equipped with the Wasserstein-$p$ distance, denoted by $W_{p}(\cdot,\cdot)$. If $\rho$ is a probability measure, $\rho^{\otimes N}$ denotes the probability space obtained by coupling $\rho$ independently $N$ times.
$\|\cdot\|_{p}$ denotes the usual $l^{p}$ vector norm in the Euclidean space, $\|\cdot\|_{L^{1}({\rho})}$ denotes $L^{1}$ norm of a function with respect to $\rho$ and $|\cdot|$ denotes the absolute value of a real number. $B^{\infty}(x,r)$ denotes the closed $l^{\infty}$ ball centered at $x$ with radius $r.$ $\mathbb{I}_{d}$ denotes the $d\times d$ identity matrix. When $u$ is a vector, $\text{diag}(u)$ denotes the diagonal matrix with $u$ being the diagonal. When $\phi$ is a function and $\mu$ is a measure, $\langle\phi,\mu\rangle$ denotes the pairing between, i.e., $\int\phi\,d\mu$. When $u$ and $v$ are vectors, $\langle u,v\rangle$ denotes the inner product in the Euclidean space.
Throughout this paper, we use the symbols $C$ and $L$ to represent generic positive uniform constants. It is important to note that these constants may take on different values in different sections or parts of this paper.
2. The dynamics of the constrained consensus-based optimization algorithm
In this section, we carry out the continuous-in-time dynamics of our method. The practical discretized algorithm will be introduced in Section 5.
Consider the following constrained optimization problem,
(1) $\min_{v\in\mathbb{R}^{d}}\ \mathcal{E}(v)\quad\text{s.t.}\quad{g(v)=0}.$
Here, we require the function $g(x)$ to be first-order differentiable. It is noteworthy that Problem (1) can be reformulated equivalently as follows:
(2) $\min_{v\in\mathbb{R}^{d}}\ \mathcal{E}(v)\quad\text{s.t.}\quad G(v)=0.$
where $G(v)={g^{2}(v)}$. Our method will be based on formulation (2).
To start with, we take $N$ particles $V^{1,N},V^{2,N},...,V^{N,N}$, which are independently sampled from a common initial law $\rho_{0}$ at initialization. Here we use $V^{{i,N}}_{t}$ for the location of the $i$-th particle at time $t$ and $d\hat{\rho}_{t}^{N}(v)=\tfrac{1}{N}\sum_{i=1}^{N}\delta_{V^{i,N}_{t}}(v)$ to denote the empirical measure. The goal of the dynamics is to encourage the measure $d\hat{\rho}_{t}^{N}$ to converge to the measure $\delta_{v^{*}}$, which is the Dirac measure at the solution of the constrained optimization problem (2). Now we propose the dynamics of the $i$-th particle, which follows the below stochastic differential equation:
(3) $\begin{gathered}dV^{i,N}_{t}=-\lambda\left(V^{i,N}_{t}-v_{\alpha}\left(\hat{\rho}_{t}^{N}\right)\right)\,dt-\dfrac{1}{\epsilon}\nabla G\left(V^{i,N}_{t}\right)\,dt+\sigma D^{i,N}_{t}\,dB^{i,N}_{t},\end{gathered}$
with $V^{i,N}_{0}\sim\rho_{0}$, where
(4) $v_{\alpha}(\hat{\rho}^{N}_{t})=\int v\cdot\dfrac{\omega_{\alpha}(v)}{\|\omega_{\alpha}\|_{L^{1}(\hat{\rho}_{t}^{N})}}\,d\hat{\rho}_{t}^{N}.$
The dynamics are driven by three distinct terms. The first and third terms are inherited from classical consensus-based optimization methods, while the second term is crafted as a forcing term to enforce the constraint. We will now explain each of them in sequence.
[…]
The choice of the consensus point is inspired by the well-known Laplace’s principle [7, 17, 38]. According to this principle, for any absolutely continuous probability measure $\rho$ on $\mathbb{R}^{d}$, one has $\lim_{\alpha\to\infty}\left(-\tfrac{1}{\alpha}\log\textstyle\int\omega_{\alpha}(v)d\rho(v)\right)=\inf_{v\in\mathrm{supp}(\rho)}\mathcal{E}(v).$ It is expected that the consensus point $v_{\alpha}(\hat{\rho}_{t}^{N})$ serves as a reasonable smooth approximation of $\operatornamewithlimits{argmin}_{i=1,...,N}\mathcal{E}(V^{i,N}_{t})$ when $\alpha$ is sufficiently large. Consequently, the particles are gathered to a location where $\mathcal{E}(v)$ attains a small value.
The diffusion term $\sigma D^{i,N}_{t}\,dB^{i,N}_{t}$ encourages particles to explore the landscape of $\mathcal{E}(v)$, where $D^{i,N}_{t}$ is a $d\times d$ matrix function that determines the way in which particles explore the landscape and $\{B_{t}^{i,N}\}_{i=1,...,N}$ are independent Wiener processes. There are different choices for the matrix function $D^{i,N}_{t},$ see [41, 13]. In this paper, we use the anisotropic exploration defined as,
(5) $D^{i,N}_{t}=\text{diag}\left(V^{i,N}_{t}-v_{\alpha}\left(\hat{\rho}_{t}^{N}\right)\right).$
It is first introduced in [13], which aims to address the curse of dimensionality.
The third term $\tfrac{1}{\epsilon}\nabla G\big(V^{i,N}_{t}\big)\,dt,$ addresses the constraint $\{G=0\}$. Since 0 is the minimum of the non-negative function $G(v)$, finding the constraint $\{G(v)=0\}$ is the same as minimizing $G(v)$. Therefore, we propose the third term as a gradient descent of $G(v)$, allowing $G(v)$ to be minimized during the algorithm’s progression. Here $\epsilon>0$ is a parameter that controls the magnitude of this term. When $\epsilon$ is small, this term will encourage particles to concentrate around $\{G=0\}$. These ideas were used in kinetic equations for swarming including alignment terms of Cucker-Smale type in order to derive kinetic models on the sphere such as the Viczek-Fokker-Planck model, see [9, 10, 11, 1] for instance. Inevitably, taking $\epsilon$ small can introduce numerical stiffness and instability, so this term must be handled carefully in the discretization. In Section 5, a stable discrete scheme is derived to treat this term.
Before we proceed to the theoretical analysis of the model, we first present a preliminary numerical result in Figures 1 and 2 to illustrate the superior performance of the proposed interacting particle system (3) compared to the projected CBO system [19] and the penalized CBO system [8] on a two-dimensional Ackley function (shown in Figure 1) with different constraints.
We defer algorithmic formulation to Section 5, and details of the preliminary experiments to Appendix K.1. Our method achieves a $100\%$ success rate in finding the unconstrained minimizer and demonstrates the fastest convergence rate in all three cases as shown in Figure 1 (b). The projected CBO performs similarly to our method when the constraint is a circle, but it is not directly applicable to parabolic curves [^1] For parabolic curve $v_{1}^{2}=v_{2}$, each evaluation involves computing the closest-point projection by solving a cubic optimality condition and selecting the minimizing real root. This per-iteration overhead is considerably higher than for the other approaches, so we omit this experiment in the comparison.. In contrast, the Penalized CBO exhibits a significantly lower success rate due to two main reasons: First, when the constrained minimizer is not a local minimizer of the objective function, the global minimizer $v_{p}^{*}$ of the penalized objective function usually differs from the constrained minimizer $v^{*}$. Second, although it is possible to increase the penalty sufficiently to reduce the distance $\|v_{p}^{*}-v^{*}\|_{2}$, the landscape is dominated by the penalized term, making the objective function resemble a minor perturbation around the penalty. Consequently, it becomes more challenging for the optimization method to locate the global minimizer, and leads to a longer time for CBO to converge.
(a) Ackley function.
(b) The success rate and averaged Euclidean distance to the constrained minimizer.
Figure 1. Objective function and success rate of three constrained CBO methods.
(a) Constrained on Sphere: The constrained minimizer is the same as the unconstrained minimizer.
(b) Constrained on Sphere: The constrained minimizer is NOT the same as the unconstrained minimizer.
(c) Constrained on parabolic curve: The constrained minimizer is NOT the same as the unconstrained minimizer.
Figure 2. The averaged distance to the true constrained minimizer over $100$ simulations.
3. Well-posedness and Mean-field limit
In this section we study some theoretical properties of the particle system described by Equation (3). We consider anisotropic diffusion (5) in both Section 3 and Section 4. Consequently, for i=1,,Ni=1,...,N, the system defined by Equation (3) transforms into the following form:
(6) $dV^{i,N}_{t}$ $=-\lambda\left(V^{i,N}_{t}-v_{\alpha}\left(\hat{\rho}_{t}^{N}\right)\right)\,dt-\dfrac{1}{\epsilon}\nabla G\left(V^{i,N}_{t}\right)\,dt$
$\quad+\sigma\,\text{diag}\left(V^{i,N}_{t}-v_{\alpha}\left(\hat{\rho}_{t}^{N}\right)\right)\,dB^{i,N}_{t},\quad V^{i,N}_{0}\sim\rho_{0},$
When the number of particles $N$ is large enough, one could study the mean-field limit as $N\rightarrow\infty$. This limit yields an equation that characterizes the macroscopic behavior of the particles, specifically their density distribution. The investigation of the mean-field equation reveals the long-term dynamics of the particle system, which is related to the convergence of the particle system or the optimization method. However, prior to this analysis in Section 4, it is necessary to establish the existence of the mean-field limit. In this section, we establish the well-posedness of Equation (6), its mean-field limit, and the well-posedness of the resultant mean-field model.
Throughout this section, we make the following assumptions.
Assumption 1.
(1) The function $\mathcal{E}$ is bounded with $\inf\mathcal{E}=\underline{\mathcal{E}}$ and $\sup\mathcal{E}=\overline{\mathcal{E}}$.
(2) There exist positive numbers $L$ and $C$ such that $\forall u,v\in\mathbb{R}^{d}$,
$\|\mathcal{E}(u)-\mathcal{E}(v)\|_{2}\leq L(\|u\|_{2}+\|v\|_{2})\|u-v\|_{2},$
$\mathcal{E}(u)-\underline{\mathcal{E}}\leq C(1+\|u\|_{2}^{2}).$
(3) There exists $L>0$ such that $\forall u,v\in\mathbb{R}^{d}$,
$\|\nabla G(u)-\nabla G(v)\|_{2}\leq L\|u-v\|_{2}.$
(4) There exists $C>0$ such that $\forall u\in\mathbb{R}^{d}$,
$\|\nabla G(u)\|_{2}\leq C\|v\|_{2}.$
Briefly speaking, in Assumption 1 (1) and (2), we assume the loss function $\mathcal{E}$ is bounded, locally Lipschitz and with at most quadratic growth. In Assumption 1 (3) and (4), we assume the gradient of the function $G$ is globally Lipschitz and with at most linear growth. Now we establish the well-posedness of the interacting particle system (6) in the following theorem.
Theorem 3.1.
(Proof in Appendix A.1) For any $N\in\mathbb{N}$, the stochastic differential equation (6) has a unique strong solution $\{V^{i,N}_{t}|t\geq 0\}_{i=1}^{N}$ for any initial condition $V^{i,N}_{0}$ satisfying $\mathbb{E}[\|V^{i,N}_{0}\|_{2}^{2}]<\infty$.
By letting the number of agents $N\rightarrow\infty$ in the model (6), the mean-field limit of the model is formally given by the following SDE
(7) $d\bar{V}_{t}=-\lambda\Big(\bar{V}_{t}-v_{\alpha}\big(\rho_{t}\big)\Big)\,dt-\dfrac{1}{\epsilon}\nabla G\,dt+\sigma\text{diag}\bigg(\bar{V}_{t}-v_{\alpha}\big(\rho_{t}\big)\bigg)\,dB_{t}.$
Then the corresponding Fokker-Planck equation is
(8) $\partial_{t}\rho_{t}=\lambda\,\text{div}\left(\left(v-v_{\alpha}\left(\rho_{t}\right)+\dfrac{1}{\epsilon}\nabla G\right)\rho_{t}\right)+\dfrac{\sigma^{2}}{2}\sum_{k=1}^{d}\partial_{x_{k}x_{k}}\left(\left(v-v_{\alpha}\left(\rho_{t}\right)\right)_{k}^{2}\rho_{t}\right).$
Next, we will prove the above equations (7), (8) are well-posed, and they model the mean-field limit. For the corresponding Fokker-Planck equation, we in particular study its weak solution, which is defined as follows.
Definition 3.2.
We say $\rho_{t}\in\mathcal{C}\left([0,T],\mathcal{P}_{4}\left(\mathbb{R}^{d}\right)\right)$ is a weak solution to (8) if
(i) It admits continuity in time in $\mathcal{C}^{{}^{\prime}}_{b}$ topology, i.e., $\left\langle\phi,\rho_{t_{n}}\right\rangle\rightarrow\left\langle\phi,\rho_{t}\right\rangle$, $\forall\phi\in\mathcal{C}_{b}\big(\mathbb{R}^{d}\big)$ and $t_{n}\rightarrow t$.
(ii) For all $\phi\in\mathcal{C}^{2}_{c}(\mathbb{R}^{d})$, it holds that
$\frac{d}{dt}\langle\phi,\rho_{t}\rangle$ $=-\lambda\langle(v-v_{\alpha}\big(\rho_{t}\big))\cdot\nabla\phi,\rho_{t}\rangle-\frac{1}{\epsilon}\langle\nabla G\cdot\nabla\phi,\rho_{t}\rangle$
$\quad+\frac{\sigma^{2}}{2}\sum_{k=1}^{d}\langle(v-v_{\alpha}\big(\rho_{t}\big))_{k}^{2}\partial_{kk}\phi,\rho_{t}\rangle,$
Remark 1.
In the Definition 3.2 (ii), the test function space is $\mathcal{C}^{2}_{c}(\mathbb{R}^{d})$. We could extend $\mathcal{C}^{2}_{c}(\mathbb{R}^{d})$ to a larger space $\mathcal{C}^{2}_{*}(\mathbb{R}^{d})$ as explained in Appendix D, which will be used in the proof later. $\mathcal{C}^{2}_{*}(\mathbb{R}^{d})$ is defined below.
$\mathcal{C}^{2}_{*}(\mathbb{R}^{d}):=\big\{$ $\phi\in\mathcal{C}^{2}(\mathbb{R}^{d})\mid|\partial_{k}\phi(x)|\leq C(1+|x_{k}|)$
$\text{ and }\sup_{x\in\mathbb{R}^{d}}|\partial_{kk}\phi(x)|<\infty\text{ for all $k=1,2,...,d$}\big\}.$
In other words, if $\rho_{t}\in\mathcal{C}\left([0,T],\mathcal{P}_{4}\left(\mathbb{R}^{d}\right)\right)$ solves equation (8) in the weak sense as in Definition 3.2, then the equality in Definition 3.2 will hold for any test function $\phi\in\mathcal{C}_{*}^{2}\big(\mathbb{R}^{d}\big)$.
Now we state the well-posedness result of (7) and (8).
Theorem 3.3.
(Proof in Appendix A.2) Let $\mathcal{E}$ satisfy Assumption 1 and $\rho_{0}\in\mathcal{P}_{4}\left(\mathbb{R}^{d}\right)$. Then given $T>0$, there exists a unique nonlinear process $\bar{V}\in\mathcal{C}\left([0,T],\mathbb{R}^{d}\right)$, satisfying (7) with initial distribution $\bar{V}_{0}\sim\rho_{0}$ in the strong sense, and $\rho_{t}=\text{Law}\left(\bar{V}_{t}\right)\in\mathcal{C}\left([0,T],\mathcal{P}_{4}\left(\mathbb{R}^{d}\right)\right)$ satisfies the corresponding Fokker-Planck equation (8) in the weak sense with $\lim_{t\rightarrow 0}\rho_{t}=\rho_{0}$.
Then we are ready to present the result showing that (7), (8) indeed characterize the mean-field limit of the particle system.
Theorem 3.4.
(Proof in Appendix A.3) Let $\mathcal{E}$ satisfy Assumption 1 and $\rho_{0}\in\mathcal{P}_{4}\left(\mathbb{R}^{d}\right)$. For any $N\geq 2$, assume that $\{(V^{i,N}_{t})\}_{i=1}^{N}$ is the unique solution to (6) with $\rho_{0}^{\otimes N}$ distributed initial data $\{V_{0}^{i,N}\}_{i=1}^{N}$. Then the limit (denoted by $\rho_{t}$) of the sequence $\{\hat{\rho}_{t}^{N}\}_{N\in\mathbb{N}}$, as $N\to\infty$ exists. Moreover, $\rho_{t}$ is deterministic and it is the unique weak solution to the corresponding Fokker-Planck equation (8) of the mean-field model.
4. Convergence to the constrained minimizer in the mean-field limit
In this section, we will analyze the behavior of the weak solution of the Fokker-Planck equation (8), which charaterizes the mean-field behavior of the proposed finite-particle system. Throughout this section, we assume Problem (2) admits a unique solution $v^{*}$, which is a common assumption in both unconstrained and constrained CBO-related literature, e.g., see [8, 13, 20, 23] [^2] In [8], the uniqueness assumption is slightly weaker: it requires the penalized loss function to admit a unique global minimizer, which coincides with the constrained minimizer.. Our primary goal is to establish a key result: under suitable assumptions and the selection of appropriate parameters, the particles will concentrate around $v^{*}$ with arbitrary closeness, confirming the effectiveness of the method in the mean-field limit.
For simplicity and without loss of generality, we assume $\mathcal{E}(v^{*})=0$. Throughout this section, we use $\rho_{t}$ to represent the solution of Equation (8) as defined in Definition 3.2, and assume it admits a density with respect to the Lebesgue measure. With a slight abuse of notation, we also use $\rho_{t}$ to refer to its density function.
4.1. Main Results
To study the convergence of $\rho_{t}$ to $v^{*}$, we define the following energy functional
(9) $\mathcal{V}\big(\rho_{t}\big):=\dfrac{1}{2}\int\|v-v^{*}\|_{2}^{2}\,d\rho_{t}(v).$
The above defined quantity $\mathcal{V}\big(\rho_{t}\big)$ provides a measure of the distance between the distribution of the particles $\rho_{t}$ and the Dirac measure at $v^{*}$, denoted as $\delta_{v^{*}}$. Specifically, we have the relationship $2\mathcal{V}\big(\rho_{t}\big)=W^{2}_{2}\big(\rho_{t},\delta_{v^{*}}\big)$, where $W_{2}\big(\rho_{t},\delta_{v^{*}}\big)$ denotes the Wasserstein-2 distance between $\rho_{t}$ and $\delta_{v^{*}}$. The diminishing behavior of $\mathcal{V}\big(\rho_{t}\big)$ indicates that $\rho_{t}$ is approaching $\delta_{v^{*}}$, implying that particles are concentrating around $v^{*}$. In this paper, we establish the following main theorem concerning the decay of $\mathcal{V}\big(\rho_{t}\big)$.
Theorem 4.1.
Suppose $G$ and $\mathcal{E}$ satisfy Assumption 2 (well-behaved). Fix any $\tau\in(0,1)$ and parameters $\lambda,\sigma>0$ with $2\lambda>\sigma^{2}$. There exists a function $I:\mathbb{R}\rightarrow\mathbb{R}$ such that for any error tolerance $\delta\in\big(0,\mathcal{V}(\rho_{0})\big)$, as long as $\rho_{0}\big(B(v^{*},r)\big)>0$ for all $r>0$ and $\int G\,d\rho_{0}(v)\leq I(\delta)$, then one can find $\alpha$ and $\epsilon$ so that
(10) $\min_{t\in[0,T^{*}]}\mathcal{V}\left(\rho_{t}\right)\leq\delta,\quad\text{where}\quad T^{*}=\dfrac{1}{(1-\tau)(2\lambda-\sigma^{2})}\log\left({\mathcal{V}\left(\rho_{0}\right)}/{\delta}\right).$
Furthermore, until $\mathcal{V}\big(\rho_{t}\big)$ reaches the prescribed accuracy $\delta$, the following exponential decay holds:
(11) $\mathcal{V}\big(\rho_{t}\big)\leq\mathcal{V}(\rho_{0})\exp\big(-(1-\tau)(2\lambda-\sigma^{2})t\big).$
Remark 2.
In the above theorem, the function $I$ only depends on $G$, $\mathcal{E}$ and parameters $\tau,\lambda,\sigma$. It does not depend on $\delta$. The choice of $\alpha,\epsilon$ will depend on $\delta$ as described in (24) and (73) respectively. In broad terms, when $\delta$ is fixed, we select a sufficiently large $\alpha$, and subsequently, based on this chosen $\alpha$, we select a small enough $\epsilon$. Additionally, it is worth noting that the selection of $\lambda$ and $\sigma$ remains independent of the dimension $d$, as the only requirement is $2\lambda>\sigma^{2}$. However, $\alpha$ will exhibit a logarithmic dependence on $d$ as illustrated in (24). Intuitively, the initial condition $\int G\,d\rho_{0}(v)\leq I(\delta)$ requires the initial particles to be near the constraint set $\{G=0\}$. This is not an unusual assumption in constrained CBO literature; for example, in [20], the convergence result assumes that the initial distribution $\rho_{0}$ is fully supported on the constraint set. While this condition is primarily technical, arising from the proof techniques, it is not essential in practice for the algorithm to converge, as demonstrated by the numerical examples in Section 5. Notably, our assumption does not require the particles to be close to the constrained minimizer $v^{*}$.
4.2. Assumption
In this subsection, we define clearly what it means by being ‘well-behaved’. $G$ and $\mathcal{E}$ are well-behaved if the following Assumption 2 is satisfied. It is worth noting that Assumption 2 in this section is independent of Assumption 1. In other words, for the proofs in this section, Assumption 1 is not required.
Assumption 2.
A. Assumptions on $\mathcal{E}$:
  1. (A1)
    $\mathcal{E}$ is bounded: $\underline{\mathcal{E}}\leq\mathcal{E}\leq\overline$$\mathcal{E}$}.
  2. (A2)
    $\mathcal{E}$ is locally Hölder continuous around $v^{*}$, i.e. there exists $r_{0}>0$ such that $\forall v_{1},v_{2}\in B^{\infty}(v^{*},r_{0})$, $|\mathcal{E}(v_{1})-\mathcal{E}(v_{2})|\leq C\|v_{1}-v_{2}\|_{\infty}^{\beta}$ for some $C\geq 0$ and $\beta>0$.
B. Assumptions on $G$:
  1. (B1)
    $\left\langle\nabla G(v),v-v^{*}\right\rangle\geq 0$ holds for any $v\in\mathbb{R}^{d}$.
  2. (B2)
    $G(v)\in\mathcal{C}^{2}_{*}\big(\mathbb{R}^{d}\big)$ and there exists $C>0$ such that $G(v)\leq C\|\nabla G(v)\|_{2}^{2}$, $\forall v\in\mathbb{R}^{d}$.
  3. (B3)
    $\nabla G(v)\neq 0$, $\forall v\in\big\{G(v)\in(0,u_{0})\big\}$ and $\int_{G(v)\in(0,u_{0})}\frac{1}{\|\nabla G(v)\|_{2}}\,dv<\infty$ for some $u_{0}>0$ small enough.
Remark 3.
Assumption 2 (B1) is related to the convexity of function $G$ but is less stringent than the convexity condition. If it is not satisfied, similar to other gradient descent algorithms, there is a possibility for some particles to get trapped in the local minimizers $\hat{v}$ of $G$, i.e. $\nabla g(\hat{v})=0$. Nevertheless, provided the function values $\mathcal{E}(v)$ at those local minimizers of $G$ do not fall below $\mathcal{E}(v^{*})$, a condition attainable by adding a positive scalar multiple of $G$ to $\mathcal{E}$ without altering the solution $v^{*}$, it will not affect the convergence of the consensus point to the constrained minimizer $v^{*}$, as evidenced in the experiments detailed in Section 5.2.1, Figure 5. It is noteworthy that this slight adjustment on $\mathcal{E}(v)$ differs from the penalization method outlined in [8]. Here, there is no necessity for the penalty parameter to approach infinity, as the convergence is enforced through the dynamics rather than penalization. The introduction of a positive scalar multiple of $G$ to $\mathcal{E}$ is to avoid the extreme case. Thus a mild penalization would suffice.
Assumption 2 (B2) is primarily technical in nature. Assumption 2 (B3) guarantees that the gradient of $G$ around the constraint $\{G=0\}$ does not vanish too rapidly.
Although Assumption 2 (B) is formulated for an abstract function $G$, it is in fact satisfied by a rather large class of functions arising from geometric constraints. The next lemma shows that Assumption 2 (B) holds for the choice where $G$ is the squared distance to a compact convex constraint set $K$ of small intrinsic dimension. This covers in particular many constrained optimization problems in which the feasible set is contained in a lower–dimensional affine subspace, while projections onto $K$ remain computable, including common combinations of linear equality/inequality constraints with norm bounds.
A practical example is portfolio optimization in finance, where $v\in\mathbb{R}^{d}$ denotes portfolio weights (with $v_{j}$ the position in asset $j$), see [40]. Typical constraints include a budget constraint $\sum_{i=1}^{d}v_{i}+c=0$, market neutrality $\beta\cdot v=0$ (with $\beta$ the vector of asset betas), diversification $\|v\|_{2}^{2}\leq D$, and leverage $\|v\|_{1}\leq 2$. Intersecting such constraints yields a compact convex feasible set whose intrinsic dimension can be at most d2d-2. Common optimization objectives in this application include, for example, variance [35], $\beta$-Value-at-Risk [33, 36], and Conditional Value-at-Risk [45].
Lemma 4.2.
[Proof in Appendix B] Let $K\subset\mathbb{R}^{d}$ be a nonempty compact convex set whose affine hull has dimension at most d2d-2. Define
$g(v):=\operatorname{dist}(v,K),\qquad G(v):=g(v)^{2},\qquad v\in\mathbb{R}^{d}.$
Then $G$ satisfies Assumption 2 (B).
C. Assumptions on the coupling of $\mathcal{E}$ and $G$: There exist a non-negative increasing function $\tau_{1}(x)$ from $\mathbb{R}$ to $\mathbb{R}$ with $\lim_{x\rightarrow 0}\tau_{1}(x)=0$, $\eta>0$, $\mu>0$, $R_{0}>0$ and $\mathcal{E}_{\infty}>0$, such that the following holds $\forall u\in[0,u_{0}]$ where $u_{0}>0$ is a small constant.
  1. (C1)
    There exist $v_{u}\in\mathbb{R}^{d},\underline{\mathcal{E}_{u}}\in\mathbb{R}$ such that $v_{u}=\mathop{\arg\min}\limits_{v\in\{G=u\}}\mathcal{E}(v)\text{ and }\underline{\mathcal{E}_{u}}=\mathcal{E}(v_{u}).$ Moreover, it holds that
    $\|v_{u}-v^{*}\|_{\infty}\leq\tau_{1}(u)\text{ and}$
    $\partial B^{\infty}\big(v_{u},r\big)\cap\big\{v\mid G(v)=u\big\}\neq\emptyset,\ \forall\text{ $r$ small}.$
  2. (C2)
    It holds that
    $\|v-v_{u}\|_{\infty}\leq\dfrac{1}{\eta}\big(\mathcal{E}(v)-\underline{\mathcal{E}_{u}}\big)^{\mu}$
    $\forall v\in B^{\infty}\big(v_{u},R_{0}\big)\cap\big\{G(v)=u\big\}$ and
    $\mathcal{E}_{\infty}<\mathcal{E}(v)-\underline{\mathcal{E}_{u}}$
    $\forall v\in B^{\infty}\big(v_{u},R_{0}\big)^{c}\cap\big\{G(v)=u\big\}$.
Remark 4.
The above Assumption 2 (C1) ensures the geometry of $\mathcal{E}$ on the set $\{G=u\}$ is similar among small enough $u$, i.e., on sets $\{G=u\}$, the constrained minimizers $v_{u}$ and constrained minimums $\mathcal{E}(v_{u})$ are close among small enough values for $u$. To illustrate, if this condition is not met, as depicted in Figure 3 (a), the desired constrained minimizer $v^{*}$ (depicted as a solid green pentagon) is considerably distant from the minimizer $v_{u}$ on a nearby level set $\{G=u\}$ (depicted as a solid orange circle) for all sufficiently small $u$. Consider an extreme case where we assume $\mathcal{E}(v_{u})$ is significantly smaller than $\mathcal{E}(v^{*})$ for all positive but sufficiently small $u$. Due to the nature of gradient descent on $G$ for each particle, which may not precisely enforce each particle to remain on the constraint $\{G=0\}$, these particles will tend to remain in a neighborhood of $\{G=0\}$. Consequently, numerous particles will cluster around $\{G=u\}$ for sufficiently small $u$, as illustrated in Figure 3 (a). Given that numerous particles are near $v_{u}$, where the function value is significantly small, the algorithm computes the consensus point around $v_{u}$ rather than $v^{*}$. Consequently, the consensus point will gradually lead particles to concentrate around $v_{u}$ rather than $v^{*}$, as illustrated in Figure 3 (b), which leads to a failure in this extreme scenario. To avoid the occurrence of such extreme cases, we proposed Assumption 2 (C1). In conjunction with other assumptions, the similarity of the local geometry of $\mathcal{E}$ on the set $\{G=u\}$ for sufficiently small values of $u$ is guaranteed as established in Lemma 4.4.
The Assumption (C2) (C2) ensures that the constrained minimizer is distinguishable from other points, i.e., on each adjacent level set $\{G=u\}$, there is a unique minimizer $v_{u}$ and $\mathcal{E}(v)$ behaves like $\|v-v_{u}\|_{\infty}^{1/\mu}$ near the $v_{u}$. This type of local coercivity condition is common in the CBO-related literature, e.g., [8, 23].
Figure 3. The blue curves represent function values on the constraint set $\{G=0\}$ and the red curves on the level set $\{G=u\}$. Dashed lines represent corresponding constraint sets. The green pentagon denotes the constrained minimizer $v^{*}$, and the orange circle represents the minimizer $v_{u}$ on the nearby level set G=u{G=u}. Empty circles represent particles.
4.3. Sketch of the Proof
In this subsection, we layout the strategy of the proof and main Lemmas, and the complete proof of Theorem 4.1.
We first plug $\tfrac{1}{2}\|v-v^{*}\|_{2}^{2}$ into Definition 3.2, which yields the following differential inequality that describes the dynamics of of the energy functional $\mathcal{V}\big(\rho_{t}\big)$.
Lemma 4.3.
(Proof in Appendix E) Let $\mathcal{V}\big(\rho_{t}\big)$ be the energy functional defined in (9). Under Assumption 2,
(12) $\dfrac{d}{dt}\mathcal{V}\big(\rho_{t}\big)\leq$ $-(2\lambda-\sigma^{2})\mathcal{V}\big(\rho_{t}\big)+\sqrt{2}(\lambda+\sigma^{2})\sqrt{\mathcal{V}\big(\rho_{t}\big)}\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{2}$
$+\dfrac{\sigma^{2}}{2}\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{2}^{2}-\dfrac{1}{\epsilon}\int\Bigl\langle\nabla G,v-v^{*}\Bigr\rangle\,d\rho_{t}(v).\vskip-6.45831pt$
It is noteworthy that if $\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{2}$ could be bounded by a suitable scalar multiple of $\sqrt{\mathcal{V}\big(\rho_{t}\big)}$, and the last term is negative, we would then obtain the inequality:
(13) $\dfrac{d}{dt}\mathcal{V}\big(\rho_{t}\big)\leq\big(-(1-\tau)(2\lambda-\sigma^{2})\big)\mathcal{V}\big(\rho_{t}\big),\vskip-4.30554pt$
to which Gronwall’s inequality can be applied, ensuring exponential decay. To this end, we first present the following auxiliary lemma, which draws out additional consequences from Assumption 2 (C).
Lemma 4.4.
(Proof in Appendix C) There exist non-negative increasing function $\tau_{2}(x),\tau_{3}(x)$ and $\tau_{4}(x)$ mapping from $\mathbb{R}$ to $\mathbb{R}$ with $\lim_{x\rightarrow 0}\tau_{i}(x)=0$ for $i=2,3,4$ so that the following hold $\forall u,r\geq 0$ small enough:
$|\underline{\mathcal{E}_{u}}|=|\mathcal{E}(v_{u})|\leq\tau_{2}(u);\quad|\mathcal{E}_{r}^{u}-\mathcal{E}_{r}|\leq\tau_{3}(\max\{u,r\});\quad|\mathcal{E}^{u}_{r}-\mathcal{E}^{0}_{r}|\leq\tau_{4}(\max\{u,r\}),$
where $\mathcal{E}^{u}_{r}=\max_{v\in B^{\infty}(v_{u},r)\cap\{G=u\}}\mathcal{E}(v)$ and $\mathcal{E}_{r}=\max_{v\in B^{\infty}(v^{*},r)}\mathcal{E}(v).$
Using the functions given in Lemma 4.4, we establish the following inequality to control $\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{2}$.
Lemma 4.5 (A Quantitative Laplace Principle).
(Proof in Appendix G) Fix $r>0$ small enough and $u>0$ small enough. $q>0$ is a constant such that $q+\mathcal{E}_{r}^{\tilde{u}}-\underline{\mathcal{E}_{\tilde{u}}}$$\mathcal{E}_{\infty}$ is true $\forall\tilde{u}\in[0,u)$. Then
(14) $\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{2}\leq$ $2\sqrt{d}\cdot\dfrac{\big(q+\mathcal{E}^{0}_{r}+\tau_{2}(u)+\tau_{4}(\max\{u,r\})\big)^{\mu}}{\eta}$
$+\dfrac{\sqrt{d}e^{-\alpha\big(q-\tau_{3}(\max\{u,r\})\big)}}{\rho_{t}\big(B^{\infty}(v^{*},r)\big)}\int_{\{G\in(0,u)\}}\|v-v_{G(v)}\|_{2}\,d\rho_{t}(v)$
$+\dfrac{\sqrt{d}e^{-\alpha\big(q-\tau_{3}(r)\big)}}{\rho_{t}\big(B^{\infty}(v^{*},r)\big)}\int_{\{G=0\}}\|v-v^{*}\|_{2}\,d\rho_{t}(v)$
$+\sqrt{d}\tau_{1}(u)+\int_{\{G(v)\geq u\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v).$
It is observed that with an appropriate choice of $q,u,r$, provided $\rho_{t}(B(v^{*},r))$ is suitably bounded from below, as proven in Lemma 4.6 in Subsection 4.4, letting $\alpha$ be sufficiently large will make the first four terms above small enough. Concerning the last term of (12), it is related to $\int G\,d\rho_{t}(v)$, which can be controlled by Lemma 4.7 in Subsection 4.4. Consequently, we can control $\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{2}$ in such a way that (13) holds, thereby ensuring exponential decay.
Remark 5.
The framework of the proof strategy was introduced in [23]. However, in our case, the Laplace principle shown in [23] is not directly applicable. There are two challenges in our proof. First, a global valley-like structure around $v^{*}$ is required to use the Laplace principle. In the constrained case, such a structure does not hold. But one can notice that, on a nearby level set $\{G=u\}$, if a comparable constrained valley-like structure exists for the corresponding constrained minimizer $v_{u}$, a constrained Laplace principle on $\{G=u\}$ can be established for small $u$. Integrating over all $u$ yields reasonable estimates on $\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|$. Secondly, there is always mass around the unconstrained global minimum. Consequently, increasing $\alpha$ would attract all particles towards the unconstrained global minimum rather than the constrained one. Therefore, one needs to find the balance between the gradient force and the tendency towards the consensus point.
Lemma 4.5 is one of the key contributions of this paper. This lemma characterizes the distribution of the particles. The particles on the constraint set $\{G=0\}$ are characterized by the first and third terms. The second and fourth terms account for the particles near the constraint, while the final term describes the particles far from the constraint. Notably, Lemma 4.5 is independent of the dynamics and is a property of the landscape. By integrating the new quantitative Laplace principle (14) with the drift towards the constrained set (as indicated by the last term in Lemma 4.3), we are able to quantify the tendency towards the constrained global minimum.
4.4. Proof of Theorem 4.1
In this subsection, we present the complete proof of Theorem 4.1. For simplicity, in the following, we assume $\tau_{1}(u)=\tau_{2}(u)=\tau_{3}(u)=\tau_{4}(u)=u$, where $\tau_{1},\ \tau_{2},\ \tau_{3},\ \tau_{4}$ are defined in Assumption 2 (C) and Lemma 4.4. We point out that the proof technique remains valid for any choice of $\tau_{i}$ that is an increasing function and converges to 0 as $u$ approaches 0.
Before the proof, we need two more lemmas on the lower bound of $\rho_{t}\big(B^{\infty}(v^{*},r)\big)$ and the dynamics of $\int G\,d\rho_{t}(v)$.
Lower bound for $\rho_{t}\big(B^{\infty}(v^{*},r)\big)$ We establish a lower bound for $\rho_{t}\big(B^{\infty}(v^{*},r)\big)$, a crucial element for our subsequent application of the Laplace principle. We first define the mollifier $\phi_{r}(v)$ as follows
(15) $\phi_{r}(v)=\left\{\begin{aligned} &\prod_{k=1}^{d}\exp\bigg({1-\dfrac{r^{2}}{r^{2}-(v-v^{*})_{k}^{2}}}\bigg),&&\text{ if $\|v-v^{*}\|_{\infty}<r$,}\\ &0,&&\text{ else.}\end{aligned}\right$$\quad$
Then we can prove the following lemma.
Lemma 4.6.
(Proof in Appendix H) Let $B=\sup_{t\in[0,T]}\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{\infty}$. Then for all $t\in[0,T]$,
$\rho_{t}\big(B^{\infty}(v^{*},r)\big)\geq\big(\int\phi_{r}\,d\rho_{0}(v)\big)e^{-at},$
where $a=2d\max\left\{\tfrac{\lambda(\sqrt{c}r+B)\sqrt{c}}{(1-c)^{2}r}+\tfrac{\sigma^{2}(cr^{2}+B^{2})(2c+1)}{(1-c)^{4}r^{2}},\tfrac{2\lambda^{2}}{(2c-1)\sigma^{2}}\right\}$ and $c\in(1/2,1)$ is some constant satisfying $(2c-1)c\geq(1-c)^{2}.$
Dynamics of $\int G\,d\rho_{t}(v)$ In Lemma 4.3, we have gained control over $\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{2}$, yet the last term in the dynamics (12) remains to be studied, which we do now.
Lemma 4.7.
(Proof in Appendix I) Assume $\sup_{t\in[0,T]}\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{2}<\infty$ and $\sup_{t\in[0,T]}\mathcal{V}\big(\rho_{t}\big)<\infty$. Then for $\epsilon>0$ small enough, $\int G\,d\rho_{t}(v)\leq\int G\,d\rho_{0}(v)$ $\forall t\in[0,T]$.
Now we are ready to prove Theorem 4.1.
Proof of Theorem 4.1.
First we use Lemma 4.3 to derive the dynamics of $\mathcal{V}\big(\rho_{t}\big)$:
$\dfrac{d}{dt}\mathcal{V}\big(\rho_{t}\big)\leq$ $-(2\lambda-\sigma^{2})\mathcal{V}\big(\rho_{t}\big)+\sqrt{2}(\lambda+\sigma^{2})\sqrt{\mathcal{V}\big(\rho_{t}\big)}\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{2}$
$+\dfrac{\sigma^{2}}{2}\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|^{2}_{2},$
where the last term on the right-hand-side of Lemma 4.3 is omitted because of its non-positivity due to Assumption 2 (B1).
Now we define $C(t)$ as $C(t)=\min\left\{\tfrac{\tau}{2}\tfrac{(2\lambda-\sigma^{2})}{\sqrt{2}(\lambda+\sigma^{2})},\sqrt{\tau\tfrac{(2\lambda-\sigma^{2})}{\sigma^{2}}}\right\}\sqrt{\mathcal{V}\big(\rho_{t}\big)}$ , and $T_{\alpha,\epsilon}$ as $T_{\alpha,\epsilon}=\sup\Big\{t\geq 0|\text{ }\mathcal{V}\big(\rho_{t^{\prime}}\big)>\delta,\big\|v_{\alpha}(\rho_{t^{\prime}}\big)-v^{*}\|_{2}\leq C(t^{\prime})\text{ for all $t^{\prime}\in[0,t]$}\Big\}.$ As long as $\|v_{\alpha}\big(\rho_{t^{\prime}}\big)-v^{*}\|_{2}\leq C(t^{\prime})$ is true, it is straightforward to verify that $\tfrac{d}{dt}\mathcal{V}\left(\rho_{t^{\prime}}\right)\leq-(1-\tau)(2\lambda-\sigma^{2})\mathcal{V}\left(\rho_{t^{\prime}}\right).$ Thus by Gronwall’s inequality, if $t\leq T_{\alpha,\epsilon}$, one has $\mathcal{V}\big(\rho_{t}\big)\leq\mathcal{V}(\rho_{0})\exp\left(-(1-\tau)(2\lambda-\sigma^{2})t\right).$
Different choices of $(\alpha,\epsilon)$ will result in different cases as follows.
Case 1 $\left({T_{\alpha,\epsilon}}\geq T^{*}\right)$.
Notice that $\mathcal{V}(\rho_{T^{*}})\leq\mathcal{V}(\rho_{0})\exp\big(-(1-\tau)(2\lambda-\sigma^{2})t\big)=\delta.$ So we have $\min_{t\in[0,T^{*}]}\mathcal{V}\big(\rho_{t}\big)\leq\delta.$ This completes the proof.
Case 2 $\left(T_{\alpha,\epsilon}<T^{*}\text{ and }\mathcal{V}\left(\rho_{T_{\alpha,\epsilon}}\right)=\delta\right)$.
In this case, it clear that $\min_{t\in[0,T^{*}]}\mathcal{V}\big(\rho_{t}\big)\leq\mathcal{V}\big(\rho_{T_{\alpha,\epsilon}}\big)=\delta$, which completes the proof.
Case 3 $\left(T_{\alpha,\epsilon}<T^{*},\ \mathcal{V}\left(\rho_{T_{\alpha,\epsilon}}\right)>\delta\text{ and }\|v_{\alpha}\left(\rho_{T_{\alpha,\epsilon}}\right)-v^{*}\|_{2}=C\left(T_{\alpha,\epsilon}\right)\right)$.
Case 3 is the only non-trivial case. We now show that suitable choices of $\alpha$ and $\epsilon$ will make Case 3 impossible.
We pick
$q=\min\left\{\tfrac{1}{4}\left(\eta\tfrac{C\left(T_{\alpha,\epsilon}\right)}{8\sqrt{d}}\right)^{1/\mu},\,\tfrac{1}{2\sqrt{d}}\mathcal{E}_{\infty}\right\},\quad r=\min\left\{\max_{s\in(0,R_{0})}\{s|\text{ }\mathcal{E}^{0}_{s}\leq\tfrac{q}{4}\},\,\tfrac{q}{4}\right\},$
$u=\min\left\{\tfrac{1}{4}\left(\eta\tfrac{C\left(T_{\alpha,\epsilon}\right)}{8\sqrt{d}}\right)^{1/\mu},\,\tfrac{q}{4},\,\tfrac{C\left(T_{\alpha,\epsilon}\right)}{4\sqrt{d}}\right\}.$
One can verify that this choice of $q,r$ and $u$ will satisfy the assumptions of Lemma 4.5, i.e., $q+\mathcal{E}_{r}^{\tilde{u}}-\underline{\mathcal{E}_{\tilde{u}}}$$\mathcal{E}_{\infty}$ (for details, please see Appendix J).
Next, in Case 3, one has $\mathcal{V}\big(\rho_{T_{\alpha,\epsilon}}\big)>\delta$. Thus
$C\left(T_{\alpha,\epsilon}\right)=$ $\min\left\{\tfrac{\tau}{2}\tfrac{(2\lambda-\sigma^{2})}{\sqrt{2}(\lambda+\sigma^{2})},\sqrt{\tau\tfrac{(2\lambda-\sigma^{2})}{d\sigma^{2}}}\right\}\sqrt{\mathcal{V}\left(\rho_{T_{\alpha,\epsilon}}\right)}$
(16) $>C_{\delta}:=\min\left\{\tfrac{\tau}{2}\tfrac{(2\lambda-\sigma^{2})}{\sqrt{2}(\lambda+\sigma^{2})},\sqrt{\tau\tfrac{(2\lambda-\sigma^{2})}{\sigma^{2}}}\right\}\sqrt{\delta}.$
Then one can see that $q$ is bounded below by $\min\left\{\tfrac{1}{4}\left(\eta\tfrac{C_{\delta}}{8\sqrt{d}}\right)^{1/\mu},\tfrac{1}{2\sqrt{d}}\mathcal{E}_{\infty}\right\}$, denoted by $q_{\delta}$. Then $r$ and $u$ are bounded by
$\min\left\{\max_{s\in(0,R_{0})}\{s|\text{ }\mathcal{E}^{0}_{s}\leq\tfrac{1}{4}q(\delta)\},\tfrac{q(\delta)}{4}\right\},\ \text{ and }\min\left\{\tfrac{1}{4}\left(\eta\tfrac{C_{\delta}}{8\sqrt{d}}\right)^{1/\mu},\tfrac{q(\delta)}{4},\tfrac{C_{\delta}}{4\sqrt{d}}\right\}$
respectively. We use $r(\delta)$ and $u(\delta)$ to denote them. We now apply Lemma 4.5 to $\rho_{T_{\alpha,\epsilon}}$ to get
$\|v_{\alpha}\big(\rho_{T_{\alpha,\epsilon}}\big)-v^{*}\|_{2}\leq 2\sqrt{d}\cdot\dfrac{\big(q+\mathcal{E}^{0}_{r}+\tau_{2}(u)+\tau_{4}(\max\{u,r\})\big)^{\mu}}{\eta}$
$+\dfrac{\sqrt{d}e^{-\alpha\big(q-\tau_{3}(\max\{u,r\})\big)}}{\rho_{T_{\alpha,\epsilon}}\big(B^{\infty}(v^{*},r)\big)}\int_{\{G\in(0,u)\}}\|v-v_{G(v)}\|_{2}\,d\rho_{T_{\alpha,\epsilon}}(v)$
$+\dfrac{\sqrt{d}e^{-\alpha\big(q-\tau_{3}(r)\big)}}{\rho_{T_{\alpha,\epsilon}}\big(B^{\infty}(v^{*},r)\big)}\int_{\{G=0\}}\|v-v^{*}\|_{2}\,d\rho_{T_{\alpha,\epsilon}}(v)+\sqrt{d}\tau_{1}(u)$
$+\int_{\{G(v)\geq u\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{T_{\alpha,\epsilon}})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{T_{\alpha,\epsilon}}(v).$
Each of the five terms on the right-hand side of the above inequality will be individually bounded.
For the first term, one can use the definition of $q,r$ and $u$ to get
(17) $2\sqrt{d}\cdot\dfrac{\left(q+\mathcal{E}^{0}_{r}+\tau_{2}(u)+\tau_{4}(\max\{u,r\})\right)^{\mu}}{\eta}$
$\leq 2\sqrt{d}\cdot\dfrac{\left(4\cdot\tfrac{1}{4}\cdot\left(\eta\tfrac{C\left(T_{\alpha,\epsilon}\right)}{8\sqrt{d}}\right)^{1/\mu}\right)^{\mu}}{\eta}=\dfrac{C\left(T_{\alpha,\epsilon}\right)}{4},$
where the inequality above is because each term in the sum on the numerator is bounded above by $\tfrac{1}{4}\cdot\left(\eta\tfrac{C\left(T_{\alpha,\epsilon}\right)}{8\sqrt{d}}\right)^{1/\mu}$ as determined by the choice of $q,r$ and $u$.
For the second term, with the chosen values of $u$ and $r$, one can first verify
(18) $q-\tau_{3}(\max\{u,r\})=q-\max\{u,r\}\geq q-\dfrac{q}{4}>\dfrac{q}{2}.$
Then
(19) $\dfrac{\sqrt{d}e^{-\alpha\big(q-\tau_{3}(\max\{u,r\})\big)}}{\rho_{T_{\alpha,\epsilon}}(B^{\infty}(v^{*},r))}\int_{\{G\in(0,u)\}}\|v-v_{G(v)}\|_{2}\,d\rho_{T_{\alpha,\epsilon}}(v)$
$\leq\dfrac{\sqrt{d}}{\int\phi_{r(\delta)}\,d\rho_{0}}\cdot e^{a(\delta)T^{*}}\cdot e^{-\alpha q/2}\Big(\int_{\{G\in(0,u)\}}\|v-v^{*}\|_{2}+\|v^{*}-v_{G(v)}\|_{2}\,d\rho_{T_{\alpha,\epsilon}}(v)\Big)$
$\leq\dfrac{\sqrt{d}}{\int\phi_{r(\delta)}\,d\rho_{0}}\cdot e^{a(\delta)T^{*}}\cdot e^{-\alpha q/2}\Big(\sqrt{2\mathcal{V}\big(\rho_{0}\big)}+\int_{\{G\in(0,u)\}}\sqrt{d}\tau_{1}(u)\,d\rho_{T_{\alpha,\epsilon}}(v)\Big)$
$\leq\dfrac{\sqrt{d}}{\int\phi_{r(\delta)}\,d\rho_{0}}\cdot e^{a(\delta)T^{*}}\cdot e^{-\alpha q/2}\Big(\sqrt{2\mathcal{V}\big(\rho_{0}\big)}+\sqrt{d}\tau_{1}(u)\Big)$
$\leq\dfrac{\sqrt{d}}{\int\phi_{r(\delta)}\,d\rho_{0}}\cdot e^{a(\delta)T^{*}}\cdot e^{-\alpha q/2}\Big(\sqrt{2\mathcal{V}\big(\rho_{0}\big)}+\mathcal{E}_{\infty}\Big),$
where $a(\delta)=2d\max\left\{\tfrac{\lambda\left(\sqrt{c}R_{0}+C(0)\right)\sqrt{c}}{(1-c)^{2}r(\delta)}+\tfrac{\sigma^{2}\left(cR_{0}^{2}+C(0)^{2}\right)(2c+1)}{(1-c)^{4}r(\delta)^{2}},\tfrac{2\lambda^{2}}{(2c-1)\sigma^{2}}\right\}$. In the first inequality above, we used (18), the fact that $T_{\alpha,\epsilon}<T^{*}$ and Lemma 4.6 with parameter $B=\sup_{t\in[0,T_{\alpha,\epsilon}]}\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{\infty}\leq\sup_{t\in[0,T_{\alpha,\epsilon}]}\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{2}\leq\sup_{t\in[0,T_{\alpha,\epsilon}]}C(t)\leq C(0)$. In the second inequality above, we used the Cauchy inequality. Also, Assumption 2 (C1) was used to deduce $\|v-v_{G(v)}\|_{2}\leq\sqrt{d}\|v-v_{G(v)}\|_{\infty}\leq\sqrt{d}\tau_{1}\big(G(v)\big)\leq\sqrt{d}\tau_{1}(u)$. In the last inequality above, we used the definition of $u$ to deduce that $u\leq\tfrac{\mathcal{E}_{\infty}}{\sqrt{d}}.$
For the third term, similarly, one has
(20) $\dfrac{\sqrt{d}e^{-\alpha\big(q-\tau_{3}(r)\big)}}{\rho_{T_{\alpha,\epsilon}}\big(B(v^{*},r)\big)}\int_{\{G=0\}}\|v-v^{*}\|_{2}\,d\rho_{T_{\alpha,\epsilon}}(v)$
$\leq\dfrac{\sqrt{d}}{\int\phi_{r(\delta)}\,d\rho_{0}}\cdot e^{a(\delta)T^{*}}\cdot e^{-\alpha q/2}\Big(\sqrt{2\mathcal{V}\big(\rho_{0}\big)}\Big).$
For the fourth term, one has
(21) $\sqrt{d}\tau_{1}(u)=\sqrt{d}u\leq\dfrac{C\big(T_{\alpha,\epsilon}\big)}{4}.$
[…]
(22) $\|v_{\alpha}(\rho_{T_{\alpha,\epsilon}})-v^{*}\|_{2}\leq$ $\dfrac{C\big(T_{\alpha,\epsilon}\big)}{2}+2\cdot\dfrac{\sqrt{d}}{\int\phi_{r(\delta)}\,d\rho_{0}}\cdot e^{a(\delta)T^{*}}\cdot e^{-\alpha q/2}\Big(\sqrt{2\mathcal{V}\big(\rho_{0}\big)}+\mathcal{E}_{\infty}\Big)$
$+\int_{\{G(v)\geq u\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{T_{\alpha,\epsilon}})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{T_{\alpha,\epsilon}}(v).$
Now we pick $\alpha$ so that
(23) $2\cdot\dfrac{\sqrt{d}}{\int\phi_{r(\delta)}\,d\rho_{0}}\cdot e^{a(\delta)T^{*}}\cdot e^{-\alpha q/2}\Big(\sqrt{2\mathcal{V}\big(\rho_{0}\big)}+\mathcal{E}_{\infty}\Big)\leq\dfrac{1}{4}C\big(T_{\alpha,\epsilon}\big).$
It turns out that if one picks $\alpha$ to be
(24) $\alpha(\delta)=\tfrac{2}{q(\delta)}\log\left(\tfrac{8\sqrt{d}\cdot e^{a(\delta)T^{*}}\cdot\left(\sqrt{2\mathcal{V}\left(\rho_{0}\right)}+\mathcal{E}_{\infty}\right)}{C_{\delta}\int\phi_{r(\delta)}\,d\rho_{0}}\right),$
then
LHS of (23)
$\leq 2\cdot\dfrac{\sqrt{d}}{\int\phi_{r(\delta)}\,d\rho_{0}}\cdot e^{a(\delta)T^{*}}\cdot e^{-\alpha q(\delta)/2}\Big(\sqrt{2\mathcal{V}\big(\rho_{0}\big)}+\mathcal{E}_{\infty}\Big)\leq\dfrac{1}{4}C_{\delta}\leq\dfrac{1}{4}C\big(T_{\alpha,\epsilon}\big),$
where in the first and third inequalities, we used the facts that $q\geq q(\delta)$ and $C\big(T_{\alpha,\epsilon}\big)>C_{\delta}.$ We remark here that $\alpha(\delta)$ is fixed once $\delta$ is fixed. With this choice of $\alpha$, we have
(25) $\|v_{\alpha}(\rho_{T_{\alpha,\epsilon}})-v^{*}\|_{2}\leq\dfrac{3}{4}C\big(T_{\alpha,\epsilon}\big)+\int_{\{G(v)\geq u\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{T_{\alpha,\epsilon}})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{T_{\alpha,\epsilon}}(v).$
Then we can go back to estimate the last term of (22). We can deduce
(26) $\int_{\{G(v)\geq u\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{T_{\alpha,\epsilon}})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{T_{\alpha,\epsilon}}(v)$
$\leq e^{\alpha(\delta)(\overline{\mathcal{E}}-\underline{\mathcal{E}})}\int_{\{G\geq u\}}\|v-v^{*}\|_{2}\,d\rho_{T_{\alpha,\epsilon}}(v)$
$\leq e^{\alpha(\delta)(\overline{\mathcal{E}}-\underline{\mathcal{E}})}\sqrt{2\mathcal{V}\big(\rho_{0}\big)}\cdot\sqrt{\rho_{T_{\alpha,\epsilon}}\big(\{G\geq u\}\big)}$
$\leq e^{\alpha(\delta)(\overline{\mathcal{E}}-\underline{\mathcal{E}})}\sqrt{2\mathcal{V}\big(\rho_{0}\big)}\cdot\dfrac{1}{\sqrt{u}}\cdot\sqrt{\int G\,d\rho_{T_{\alpha,\epsilon}}(v)}$
$\leq e^{\alpha(\delta)(\overline{\mathcal{E}}-\underline{\mathcal{E}})}\sqrt{2\mathcal{V}\big(\rho_{0}\big)}\cdot\dfrac{1}{\sqrt{u(\delta)}}\cdot\sqrt{\int G\,d\rho_{T_{\alpha,\epsilon}}(v)},$
where in the second inequality, we used the Cauchy inequality, in the third inequality, we used the Markov inequality and in the last inequality, we used the fact that $u\geq u(\delta).$ Thus by applying Lemma 4.7 with $B=C(0)$ and $\tilde{B}=\mathcal{V}(\rho_{0})$, when $\epsilon$ is small enough (a quantification is given in Appendix I, Equation (73)), the following holds:
(27) $\int G\,d\rho_{T_{\alpha,\epsilon}}(v)\leq\int G\,d\rho_{0}(v).$
Thus combining (26) and (27) gives
(28) $\int_{\{G(v)\geq u\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{T_{\alpha,\epsilon}})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{T_{\alpha,\epsilon}}(v)$
$\leq e^{\alpha(\delta)(\overline{\mathcal{E}}-\underline{\mathcal{E}})}\sqrt{2\mathcal{V}\big(\rho_{0}\big)}\cdot\dfrac{1}{\sqrt{u(\delta)}}\cdot\sqrt{\int G\,d\rho_{0}(v)}.$
Now we pick the function $I(x)$ to be $\tfrac{1}{128\mathcal{V}\left(\rho_{0}\right)}C^{2}_{x}e^{-2\alpha(x)(\overline{\mathcal{E}}-\underline{\mathcal{E}})}u^{2}(x),$ where $C_{x}$ and $\alpha(x)$ are defined in (16) and (24) respectively. As long as
(29) $\int G\,d\rho_{0}(v)\leq I(\delta),$
combining (28) and (29) yield
$\int_{\{G(v)\geq u\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{T_{\alpha,\epsilon}})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{T_{\alpha,\epsilon}}(v)$
$\leq e^{\alpha(\delta)(\overline{\mathcal{E}}-\underline{\mathcal{E}})}\sqrt{2\mathcal{V}\big(\rho_{0}\big)}\cdot\dfrac{1}{\sqrt{u(\delta)}}\cdot\sqrt{\int G\,d\rho_{0}(v)}$
$\leq\dfrac{1}{8}C_{\delta}\leq\dfrac{1}{8}C\big(T_{\alpha,\epsilon}\big).$
By plugging the above inequality back to (25), one gets $\|v_{\alpha}(\rho_{T_{\alpha,\epsilon}})-v^{*}\|_{2}\leq\tfrac{3}{4}C\left(T_{\alpha,\epsilon}\right)+\tfrac{1}{8}C\left(T_{\alpha,\epsilon}\right)=\tfrac{7}{8}C\left(T_{\alpha,\epsilon}\right)<C\left(T_{\alpha,\epsilon}\right),$ which contradicts with the assumption $\|v_{\alpha}(T_{\alpha,\epsilon})-v^{*}\|_{2}=C\left(T_{\alpha,\epsilon}\right)$ as stated in Case 3. Therefore, we have demonstrated that under the assumptions of Theorem 4.1, if one selects $\alpha$ to be $\alpha(\delta)$ and chooses $\epsilon$ to be sufficiently small, Case 3 will not occur.
Thus we have proved that if all the conditions in Theorem 4.1 are satisfied, the desired decay can be achieved with the specified choices of $\alpha$ and $\epsilon$. ∎
5. Numerical Experiments
In this section, we present the discretized algorithm of the continuous model (3). Throughout this section, we use the anisotropic version (5) as it is more efficient in solving high-dimensional optimization problems.
5.1. Algorithm
First, one notices that in the time-continuous model (3), the forcing term $\frac{1}{\epsilon}\nabla G$ needs to be relatively large for the particles to remain near the constraint set. However, a straightforward explicit scheme of the dynamics requires the time step $\gamma$ to be of the same order as $\epsilon$. This implies that as $\epsilon$ approaches zero, the algorithm becomes expensive. On the other hand, making the stiff term $\frac{1}{\epsilon}\nabla G(V^{i}_{k+1})$ implicit enhances numerical stability, but it becomes computationally challenging for complex constraints. To address this, we introduce an algorithm with better stability for any equality constraints.
The key idea is to employ Taylor expansion to approximate the term $\nabla G(V^{i}_{k+1})$ in the implicit algorithm with its first-order approximation.
$V^{j}_{k+1}=$ $V^{j}_{k}-\lambda\gamma(V^{j}_{k}-v_{\alpha}(\hat{\rho}_{k}))$
$-\left(\frac{\gamma}{\epsilon}\nabla G(V^{j}_{k})+\frac{\gamma}{\epsilon}\nabla^{2}G(V^{j}_{k})(V_{k+1}^{j}-V^{j}_{k})\right)-\sigma\sqrt{\gamma}(V^{j}_{k}-\bar{v}_{k})\odot z_{k},$
which leads to the following constrained CBO algorithm,
(30) $V^{j}_{k+1}=V^{j}_{k}-\left[I+\frac{\gamma}{\epsilon}\nabla^{2}G(V^{j}_{k})\right]^{-1}\Bigg($ $\lambda\gamma(V^{j}_{k}-v_{\alpha}(\hat{\rho}_{k}))$
$+\frac{\gamma}{\epsilon}\sum_{i=1}^{m}g_{i}(V^{j}_{k})\nabla g_{i}(V^{j}_{k})+\sigma\sqrt{\gamma}(V^{j}_{k}-\bar{v}_{k})\odot z_{k}\Bigg),$
where $\nabla^{2}G(v)$ represents the Hessian of $G(v)$, i.e., $\nabla^{2}G(v)=\sum_{i=1}^{m}(\nabla g_{i})^{\top}\nabla g_{i}+g_{i}\nabla^{2}g_{i}$ and $\gamma$ is the time step, and $\odot$ is a point-wise multiplication, i.e., the i-th component of $x\odot y$ is $x_{i}y_{i}$. Here $V^{j}_{k}$ approximates the space location of the $j$-th particle at time $t=k\gamma$, and $z_{k}$ is a $d$-dimensional random variable following a standard normal distribution $\mathcal{N}(0,\mathbb{I}_{d})$. During different steps, $z_{k}$ is sampled independently. The complete algorithm is formulated as in Algorithm 1.
The preliminary results shown in Figure 2 (a) are obtained using the above scheme with $\epsilon=0.01$ and $\gamma=0.1$, which demonstrates the stability of the algorithm.
We also propose an alternative algorithm when the dimensionality is high, where we introduce independent noise after the particles concentrate. This version introduces additional noises to help the particles explore the landscape better, which is necessary when the dimension of the optimization problem is high. The complete algorithm is formulated as in Algorithm 2.
Algorithm 1 Constrained CBO Algorithm
Initialization: Choose hyperparameters $\epsilon$, $\alpha$, time step $\gamma$, stopping threshold $\epsilon_{\text{stop}}$, and sample size $N$. Sample $N$ particles $V^{j}$ from distribution $\rho_{0}(v)$.
1:while $\frac{1}{dN}\sum_{j=1}^{N}\|V^{j}-v_{\alpha}(\hat{\rho})\|^{2}>\epsilon_{\text{stop}}$ do
[…]
$v_{\alpha}(\hat{\rho})=\frac{1}{Z}\sum_{j=1}^{N}\mu_{j}V^{j},\quad\text{with}\quad Z=\sum_{j=1}^{N}e^{-\alpha\mathcal{E}(V^{j})},\quad\mu_{j}=e^{-\alpha\mathcal{E}(V^{j})}$
3:  Update each particle’s position $\{V^{j}\}_{j=1}^{N}$:
$\displaystyle V^{j}\leftarrow V^{j}-\left[I+\frac{\gamma}{\epsilon}\sum_{i=1}^{m}\nabla^{2}\left[g^{2}_{i}(V^{j})\right]\right]^{-1}\Bigg($ $\lambda\gamma(V^{j}-v_{\alpha}(\hat{\rho}))+\frac{\gamma}{\epsilon}\sum_{i=1}^{m}\nabla\left[g_{i}^{2}(V^{j})\right]$
$+\sigma\sqrt{\gamma}(V^{j}-v_{\alpha}(\hat{\rho}))\odot z\Bigg),$
where $z\sim\mathcal{N}(0,\mathbb{I}_{d})$.
4:end while
5:Output $)v_{\alpha}\big(\hat{\rho}\big),\mathcal{E}(v_{\alpha}\big(\hat{\rho}\big$))
Algorithm 2 Constrained CBO Algorithm with Independent Noise
Initialization: Choose suitable hyper-parameters $\epsilon,\alpha$, and time step $\gamma$, stopping threshold $\epsilon_{\text{stop}},\epsilon_{\text{indep}}$, independent noise $\sigma_{\text{indep}}$. Sample $N$ particles $V^{j}$ following distribution $\rho_{0}(v)$ and set $\mathcal{E}^{\star}$ to be a large constant.
1:while $|\mathcal{E}(v_{\alpha}\big(\hat{\rho}\big))-\mathcal{E}^{\star}|\geq\epsilon_{\text{indep}}$ do
2:  while $\frac{1}{dN}\sum_{j=1}^{N}\|V^{j}-v_{\alpha}(\hat{\rho})\|^{2}>\epsilon_{\text{stop}}$ do
[…]
$v_{\alpha}(\hat{\rho})=\frac{1}{Z}\sum_{j=1}^{N}\mu_{j}V^{j},\quad\text{with}\quad Z=\sum_{j=1}^{N}e^{-\alpha\mathcal{E}(V^{j})},\quad\mu_{j}=e^{-\alpha\mathcal{E}(V^{j})}$
4:    Update each particle’s position $\{V^{j}\}_{j=1}^{N}$:
$\displaystyle V^{j}\leftarrow V^{j}-\left[I+\frac{\gamma}{\epsilon}\sum_{i=1}^{m}\nabla^{2}\left[g^{2}_{i}(V^{j})\right]\right]^{-1}\Bigg($ $\lambda\gamma(V^{j}-v_{\alpha}(\hat{\rho}))+\frac{\gamma}{\epsilon}\sum_{i=1}^{m}\nabla\left[g_{i}^{2}(V^{j})\right]$
$+\sigma\sqrt{\gamma}(V^{j}-v_{\alpha}(\hat{\rho}))\odot z\Bigg),$
where $z\sim\mathcal{N}(0,\mathbb{I}_{d})$.
5:  end while
[…]
7:
[…]
8:  end if
9:  Each particle does an independent move:
$V^{j}\leftarrow V^{j}+\sigma_{\text{indep}}\sqrt{\gamma}z,\quad\text{for }1\leq j\leq N,$
where $z\sim\mathcal{N}(0,\mathbb{I}_{d})$.
10:end while
[…]
5.2. Numerical examples
5.2.1. A simple example
We first test the proposed algorithm on a simple objective function,
(31) $\min_{v\in\mathbb{R}^{d}}$ $\|v\|_{2}^{2}.$
We consider three types of constraints. The first is a line-segment constraint in $\mathbb{R}^{3}$ ($d=3$), with
(32) $g(v)=\mathrm{dist}(v,L),$
where $\mathrm{dist}(\cdot,L)$ denotes the Euclidean distance to the line segment $L$ with endpoints $(1.6,0.2,0.4)$ and (0.3,0.7,0.5)(-0.3,-0.7,0.5). The second case is an ellipse, with
(33) $g(v)=\frac{(v_{1}+1)^{2}}{2}+v_{2}^{2}-1=0.$
The third case is a line, with
(34) $g(v)=v_{1}+v_{2}-3=0.$
The exact minimizers of those three cases are,
${\text{Line-segment: }v^{*}=(0.2361,-0.446,0.4718)};$
$\text{Ellipse: }v^{*}=(\sqrt{2}-1,0);\quad\text{Line: }v^{*}=(3/2,3/2).$
We use Algorithm 1 with $N=50,\ \alpha=50,\ \epsilon=0.01,\ \lambda=1,\ \sigma=5,\ \gamma=0.1,\ \epsilon_{\text{stop}}=10^{-14},$ and the particles are initially set to follow a uniform distribution in $3,3]^{d}$. We consider our search for the constrained minimizer successful if, when the algorithm finishes, $\|v_{\alpha}\big(\hat{\rho}\big)-v^{*}\|_{\infty}\leq 0.1$. The success rate and the average distance are shown in Table 1, where the average distance to $v^{*}$ in the table is measured using the following norm
(35) $D(v,v^{*})=\dfrac{1}{\sqrt{d}}\left\lVert v-v^{*}\right\rVert=\left(\dfrac{1}{d}\sum_{i=1}^{d}(v-v^{*})_{i}^{2}\right)^{1/2}.$
In Figure 4, we show the evolution of the objective function value $)\mathcal{E}(v_{\alpha}\big(\hat{\rho}\big$)), the constraint value $)g(v_{\alpha}\big(\hat{\rho}\big$)), and the distance $D(v_{\alpha}\big(\hat{\rho}\big),v^{*})$ over $100$ simulations. It is evident that the consensus point stabilizes within 10 steps for all simulations.
In Figure 5, the evolution of all the particles and the consensus point are shown in time steps $k=0,5,50,100$. In all cases, after $5$ steps, most of the particles are driven to the constraints by the strong constraint term $\frac{1}{\epsilon}\nabla G(v)$ and stay there consistently. It is worth noting that, in contrast to the line-segment case, which satisfies Assumption 2 (B) (guaranteed by Lemma 4.2), in the ellipse case not all particles converge to the consensus point. Instead, some particles remain trapped at a point $\tilde{v}$ satisfying $\nabla g(\tilde{v})=0$, rather than reaching the feasible set $\{v:g(v)=0\}$. This behavior occurs when $G(v)=g^{2}(v)$ fails to satisfy Assumption 2 (B), in particular condition (B1). Nevertheless, it does not affect the convergence of the consensus point as long as the loss value at $\tilde{v}$ is not too small compared to the constrained minimum (see Remark 3).
Table 1. The result of Algorithm 1 on (31) with constraints (32), (33), and (34)
success rate average distance to $v^{*}$
line-segment constraint $100\%$ $0.0182$
ellipse constraint $100\%$ $0.0147$
line constraint $100\%$ $0.0157$
(a) blue lines: $)\mathcal{E}(v_{\alpha}\big(\hat{\rho}\big$)), red lines: $)g(v_{\alpha}\big(\hat{\rho}\big$))
(b) $D(v_{\alpha}\big(\hat{\rho}\big),v^{*})$
(c) blue lines: $)\mathcal{E}(v_{\alpha}\big(\hat{\rho}\big$)), red lines: $)g(v_{\alpha}\big(\hat{\rho}\big$))
(d) $D(v_{\alpha}\big(\hat{\rho}\big),v^{*})$
(e) blue lines: $)\mathcal{E}(v_{\alpha}\big(\hat{\rho}\big$)), red lines: $)g(v_{\alpha}\big(\hat{\rho}\big$))
(f) $D(v_{\alpha}\big(\hat{\rho}\big),v^{*})$
Figure 4. The first row is the result for the optimization problem (31) with line-segment constraint (32), the second row is the result for the ellipse constraint (33), and the third row is for the line constraint (34). The left column is the evolution of the objective function value and constraint value, while the right column is the evolution of the distance between the consensus point and the exact minimizer, where the distance is defined in (35). The light lines are results from $100$ simulations, while the dark lines are the average values.
(a) k=0
(b) k=5
(c) k=50
(d) k=100
(e) k=0
(f) k=5
(g) k=50
(h) k=100
(i) k=0
(j) k=5
(k) k=50
(l) k=100
Figure 5. Evolution of all particles (blue) and the corresponding consensus point (red) for the constrained problems: (32) with the line-segment constraint (first two rows), (31) with the ellipse constraint (33) (third row), and (31) with the line constraint (34) (fourth row). The constraint set is shown in black. The black marker denotes the unconstrained global minimizer of the objective function, and the green marker denotes the constrained minimizer $v^{*}$.
5.2.2. Ackley function
We now test the proposed algorithms on a highly non-convex objective function. Consider the following Ackley function,
(36) $\min_{v}\quad-A\exp\left(-a\sqrt{\frac{b^{2}}{d}\left\lVert v-\hat{v}\right\rVert_{2}^{2}}\right)-\exp\left(\frac{1}{d}\sum_{i=1}^{d}\cos(2\pi b(v_{i}-\hat{v}_{i}))\right)+e^{1}+A,$
where $b=1,A=20,a=0.1$, and $\hat{v}$ is the global minimum of the unconstrained problem. The above function in two-dimension is shown in Figure 1. Here we consider five different constraints,
[…]
(37) $(0.2,0.5,0.7)$ and $(0.5,0.2,0.5)$
[…]
(38) $\{v\mid v_{1}^{2}+\cdots+v_{18}^{2}\leq 0.25,\ v_{19}=v_{20}=0\}$
(39) $\text{Case 3.}\quad\|v\|_{2}^{2}-1=0\ \text{in $\mathbb{R}^{3}$ and in $\mathbb{R}^{20}$}$
(40) $\text{Case 4.}\quad\sum_{i=1}^{d-1}v_{i}^{2}-v_{d}=0\ \text{in $\mathbb{R}^{3}$ and in $\mathbb{R}^{20}$}$
(41) $\text{Case 5.}\quad g(v)=\sum_{i=1}^{d}v_{i}-1=0,\quad\displaystyle 2\sum_{i=1}^{d-1}v_{i}-\frac{1}{2}v_{d}-\frac{1}{2}=0\ \text{in $\mathbb{R}^{3}$}$
By by Lemma 4.2, Case 1 and Case 2 satisfy Assumption 2 (B). Here, we set $\hat{v}=(0.4,\cdots,0.4)$ so that the unconstrained minimizer is not the same as the constrained minimizer. The constrained minimizers for the three-dimensional cases are
${\text{Case 1. }v^{*}=(0.4003,0.2997;0.5665);}\quad\text{Case 3. }v^{*}=1/\sqrt{3}(1,1,1);$
$\text{Case 4. }v^{*}=(0.4283,0.4283,0.3669);\quad\text{Case 5. }v^{*}=(0.2,0.2,0.6).$
The constrained minimizers for the 20-dimensional case are
${\text{Case 2. }v_{i}^{*}=0.1179,\ 1\leq i\leq 18,\ v_{19}^{*}=v_{20}^{*}=0};$
$\text{Case 3. }v_{i}^{*}=1/\sqrt{20},\ 1\leq i\leq 20;$
$\text{Case 4. }v_{i}^{*}=0.3542,\ 1\leq i\leq 19,\ v_{20}^{*}=2.3839.$
For the $3$-dimensional Ackley function, we use Algorithm 1 with
$N=100,\ \alpha=50,\ \epsilon=0.01,\ \lambda=1,\ \sigma=1,\ \gamma=0.1,\ \epsilon_{\text{stop}}=10^{-14}.$
For the $20$-dimensional Ackley function, we use Algorithm 2 with
[…]
${\text{Case 2.}\ \epsilon_{\min}=0.05,\ \sigma_{\mathrm{indep}}=0.3;}$
[…]
and all the particles initially follow $V^{j}\simUnif$$3,3]^{d}$.
The evolution of the distance $D(v_{\alpha}\big(\hat{\rho}\big),v^{*})$ between the consensus point and the accurate solution is shown in Figure 6, where one can see that the consensus point converges to the true minimizer within $100$ steps. Besides, the success rate, averaged distance for the output consensus point $v^{*}$, and the averaged total steps are reported in Table 2. We consider the simulation to be successful if $\max_{k}|v_{\alpha}\big(\hat{\rho}\big)_{k}-v^{*}_{k}|\leq 0.1$, and the distance to $v^{*}$ is measured in $D(v_{\alpha}\big(\hat{\rho}\big),v^{*})$ and averaged over $100$ simulations. One can see that, except for the $20$-dimensional Cases 2 and 4, the algorithm finds the exact minimizer with a $100\%$ success rate. Moreover, except for the $3$-dimensional Case 1 and the $20$-dimensional Case 4, the algorithm achieves high accuracy within $400$ steps. Even for the $20$-dimensional Cases 2 and 4, although the success rate is slightly below $100\%$, the average distance to $v^{*}$ is below $0.05$, indicating that the iterates remain relatively close to the exact minimizer $v^{*}$.
The reason for the non-smoothness in the later stage of the average line is due to the limited number of samples for the larger steps. In most simulations, the algorithm typically concludes its iterations around the average total steps in the table. As it is difficult to obtain the exact constrained minimizer of the $20$-dimensional Ackley function under the constraints in (41), we only consider Case 5 in $\mathbb{R}^{3}$. Moreover, for the $20$-dimensional Case 2, $v^{*}$ is a reference solution computed by first projecting the unconstrained global minimizer onto $S$ and then using the MATLAB Optimization Toolbox to refine the solution, since the exact constrained minimizer is hard to obtain.
Table 2. The result of Algorithm 1 on 33-dimensional Ackley function and Algorithm 2 for 2020-dimensional Ackley function.
success rate average distance to $v^{*}$ average total steps
case 1 d=3 $100\%$ $3.8\times 10^{-3}$ 1000
case 2 d=20 $98\%$ $2.4\times 10^{-2}$ 158
case 3 d=3 $100\%$ $8\times 10^{-3}$ 295
d=20 $100\%$ $1.56\times 10^{-2}$ 390
case 4 d=3 $100\%$ $4.5\times 10^{-3}$ 213
d=20 $96\%$ $3.13\times 10^{-2}$ 4288
case 5 d=3 $100\%$ $2.8\times 10^{-3}$ 163
(a) $d=3$ case 1
(b) $d=3$ case 3
(c) $d=3$ case 4
(d) $d=3$ case 5
(e) $d=20$ case 2
(f) $d=20$ case 3
(g) $d=20$ case 4
Figure 6. The evolution of the distance $D(v_{\alpha}\big(\hat{\rho}\big),v^{*})$ between the consensus point and the exact minimizer. The objective function is the Ackley fuction (36) and the constraints are (37) - (41). The light lines are the results from $100$ simulations, while the dark lines are the average values.
5.2.3. Thomson’s Problem
The Thomson problem involves determining the positions for $k$ electrons on a sphere in a way that minimizes the electrostatic interaction energy between each pair of electrons with equal charges. The associated constrained optimization problem is formulated as follows,
$\min\quad\mathcal{E}(v_{1},.,v_{k})=\frac{1}{k}\sum_{i<j}\frac{1}{\left\lVert v_{i}-v_{j}\right\rVert_{2}}$
$\text{s.t.}\quad\left\lVert v_{i}\right\rVert^{2}_{2}-1=0,\quad\text{for }i=1,\cdots,k.$
We use Algorithm 2 with
[…]
and all the particle initially follow $V^{j}\sim\text{Unif}[-1,1]^{3k}$.
[…]
(42) $\text{relative error}=\frac{|\mathcal{E}(v_{\alpha}\big(\hat{\rho}\big))-\mathcal{E}(v^{*})|}{\mathcal{E}(v^{*})}$
[…]
$\text{relative error}\leq 0.05,\quad\sum_{i=1}^{k}(|\left\lVert v_{i}\right\rVert^{2}_{2}-1|)\leq 10^{-3}.$
In Figure 7, the evolution of the relative error across 100 simulations and their average values are depicted, illustrating that all experiments converge to the optimal minimizer within 2000 steps. The nonsmoothness of the average lines is due to the fewer samples in large steps. For $k=56,470$, corresponding to an optimization problem of dimensions $168$ and $1410$ with $56,470$ constraints, the success rate is not $100\%$. However, it remains above $90\%$. Besides, the relative error and constraints value in the third and fourth columns of Table 3 are over the success simulations, which are very small. This verifies our algorithm has an excellent performance in high dimensions.
Table 3. The result of Algorithm 2 on Thomson problem.
success rate relative error constraints value total steps
$k=2,(d=6)$ $100\%$ $4.4\times 10^{-3}$ $3.8\times 10^{-11}$ $382$
$k=3,(d=9)$ $100\%$ $9.9\times 10^{-3}$ $1.4\times 10^{-10}$ $407$
$k=8,(d=24)$ $100\%$ $1.78\times 10^{-2}$ $2.3\times 10^{-10}$ $567$
$k=15,(d=45)$ $100\%$ $1.57\times 10^{-2}$ $3.4\times 10^{-10}$ $895$
$k=56,(d=168)$ $97\%$ $1.44\times 10^{-2}$ $2.91\times 10^{-6}$ $1610$
$k=470,(d=1410)$ $93\%$ $1.95\times 10^{-2}$ $4.03\times 10^{-6}$ $1960$
(a) k=2
(b) k=3
(c) k=8
(d) k=15
(e) k=56
(f) k=470
Figure 7. Thomson Problem: the decay of the relative error over $100$ simulation and its mean.
6. Conclusions
In this paper, we propose a new CBO-based method for solving constrained non-convex minimization problem with equality constraints and potentially non-differentiable loss functions. Specifically, we augment the original CBO framework with a new forcing term designed to guide particles toward the constraint set. On the theoretical side, we conduct a rigorous analysis of the mean-field limit for the proposed model (6), deriving the corresponding macroscopic model (7) and establishing well-posedness results for both the microscopic and macroscopic models. To demonstrate the convergence of the method, we study the long-time behavior of the macroscopic model (7) through an analysis of the associated Fokker-Planck equation (8). Our results establish that, under Assumption 2 and with a proper choice of parameters, particles converge to the constrained minimizer $v^{*}$ with arbitrary closeness. Notably, Assumption 2 (C) fits well with the basic nature of the algorithm, while Assumption 2 (B) serves as a technical requirement needed by our proof technique, which might be relaxed further with an alternative proof technique, as suggested by the performance exhibited in numerical experiments where Assumption 2 (B) may not be strictly satisfied. On the practical side, we proposed a stable algorithm based on the continuous-in-time model. In Section 5, the algorithm’s performance is illustrated through a series of experiments, including challenging high-dimensional problems.
Appendix
Appendix A Some details in the Proofs of Well-posedness and Mean-field limit
A.1. Proof of Theorem 3.1
Consider the microscopic model, which is governed by the following equation:
$dV^{i,N}_{t}=$ $-\lambda\left(V^{i,N}_{t}-v_{\alpha}(\hat{\rho}_{t}^{N})\right)\,dt-\dfrac{1}{\epsilon}\nabla G\left(V^{i,N}_{t}\right)\,dt$
$+\sigma\text{diag}\left(V^{i,N}_{t}-v_{\alpha}(\hat{\rho}_{t}^{N})\right)\,dB^{i,N}_{t},$
with initial distribution $V^{i,N}_{0}\sim\rho_{0}$, where $i=1,\dots,N$. We can concantenate $\left\{V_{t}^{i,N}\right\}_{i=1}^{N}$ into one vector and put them in one equation. To be specific, we define
$V_{t}=\left(\left(V_{t}^{1,N}\right)^{T},...,\left(V_{t}^{N,N}\right)^{T}\right)^{T}.$
Then $V_{t}$ is a vector in $\mathbb{R}^{Nd}$ for each fixed $t$ and it will satisfy the following equation:
(43) $dV_{t}=-\lambda F_{N}(V_{t})\,dt-\dfrac{1}{\epsilon}L_{N}(V_{t})\,dt+\sigma M_{N}(V_{t})\,dB_{t}^{(N)}.$
Here $B^{(N)}$ is the standard Wiener process in $\mathbb{R}^{Nd}$,
$L_{N}(V_{t})=\left(\left(\nabla G\left(V_{t}^{1,N}\right)\right)^{T},...,\left(\nabla G\left(V_{t}^{N,N}\right)\right)^{T}\right)^{T}\in\mathbb{R}^{Nd},$
$M_{N}(V_{t})=\text{diag}\left(F_{N}^{1}\left(V_{t}\right),...,F_{N}^{N}\left(V_{t}\right)\right)\in\mathbb{R}^{Nd\times Nd}$
and
$F_{N}(V_{t})=\left(\left(F_{N}^{1}\left(V_{t}\right)\right)^{T},...,\left(F_{N}^{N}\left(V_{t}\right)\right)^{T}\right)^{T}\in\mathbb{R}^{Nd},$
where
$F_{N}^{i}(V_{t})=\dfrac{\sum_{j\neq i}\left(V_{t}^{i,N}-V_{t}^{j,N}\right)\omega_{\alpha}\left(V_{t}^{j,N}\right)}{\sum_{j}\omega_{\alpha}(V_{t}^{j,N})}\in\mathbb{R}^{d}.$
Thus it suffices to prove the well-posedness result of equation (43). The below theorem gives the well-posedness.
Theorem A.1.
For each $n\in\mathbb{N}$, the stochastic differential equation (43) has a unique strong solution $\{V_{t}|t\geq 0\}$ for any initial condition $V_{0}$ satisfying $\mathbb{E}[\big\|V_{0}\big\|^{2}]<\infty$.
Proof.
Following the same steps in Theorem 2.1 [12], we obtain
$-2\lambda V\cdot F_{N}(V)\leq 2\lambda\sqrt{N}\big\|V\big\|^{2},\text{ and }\text{trace}(M_{N}M_{N}^{T})(V)=\big\|F_{N}(V)\big\|^{2}\leq 4N\big\|V\big\|^{2}.$
Thus
$-2\lambda V\cdot F_{N}(V)+\sigma^{2}\text{trace}(M_{N}M_{N}^{T})(V)\leq b_{N}\big\|V\big\|^{2},$
where $b_{N}$ is a positive number that depends only on $\lambda,\sigma,d$ and $N$. Also, we notice that for $X\in\mathbb{R}^{d}$
$-\dfrac{2}{\epsilon}X\cdot\nabla G(X)\leq\dfrac{2}{\epsilon}\big\|X\big\|\cdot\big\|\nabla G(X)\big\|\leq\dfrac{2}{\epsilon}\big\|X\big\|\cdot\big\|X\big\|=\dfrac{2}{\epsilon}\big\|X\big\|^{2},$
where we used Assumption 1 (4). Thus,
$-\dfrac{2}{\epsilon}V\cdot L_{N}(V)\leq\dfrac{2}{\epsilon}\big\|V\big\|^{2}.$
This implies that
$2V\cdot\left(-\lambda F_{N}(V)-\dfrac{1}{\epsilon}L_{N}(V)\right)+\sigma^{2}\text{trace}(M_{N}M_{N}^{T})(V)\leq\tilde{b}_{N}\big\|V\big\|^{2},$
where $\tilde{b}_{N}$ is some positive number that depends only on $\lambda,\sigma,d,\epsilon$ and $N$. Then we apply Theorem 3.1 in [18] to finish the proof. ∎
A.2. Proof of Theorem 3.3
Below is Lemma 3.2 from [12].
Lemma A.2.
Let $\mathcal{E}$ satisfy Assumption 1 and $\mu,\hat{\mu}\in\mathcal{P}_{2}(\mathbb{R}^{d})$ with
$\int\big\|v\big\|^{4}\,d\mu,\int\big\|\hat{v}\big\|\,d\hat{\mu}\leq K.$
Then the following stability estimate holds
[…]
Also, we recall Theorem 11.3 in [26].
Theorem A.3.
Let $T$ be a compact mapping of a Banach space $\mathcal{B}$ into itself, and suppose there exists a constant $M$ such that
$\big\|x\big\|_{\mathcal{B}}<M$
for all $x\in\mathcal{B}$ and $\sigma\in[0,1]$ satisfying $x=\sigma Tx$. Then $T$ has a fixed point.
Proof of Theorem 3.3.
Step 1 (construct a map $T$):
Let us fix $u_{t}\in\mathcal{C}[0,T]$. By Theorem 6.2.2 in [3], there is a unique solution to
(44) $\begin{gathered}dV_{t}=-\lambda(V_{t}-u_{t})\,dt-\dfrac{1}{\epsilon}\nabla G\,dt+\sigma\text{diag}\left(V_{t}-u_{t}\right)\,dB_{t},\\ V_{0}\sim\rho_{0},\end{gathered}$
We use $\rho_{t}$ to denote the corresponding law of the unique solution. Using $\rho_{t}$, one can compute $v_{\alpha}\left(\rho_{t}\right)$, which is uniquely determined by $u_{t}$ and is in $\mathcal{C}[0,T]$. Then one can construct a map from $\mathcal{C}[0,T]$ to $\mathcal{C}[0,T]$ which maps $u_{t}$ to $v_{\alpha}\left(\rho_{t}\right)$.
Step 2 ($T$ is compact):
Firstly, from Chapter 7 in [3], we obtain the following inequality for the solution $V_{t}$ to equation (44):
$\mathbb{E}\left[\big\|V_{t}\big\|\right]^{4}\leq\left(1+\mathbb{E}\left[\big\|V_{0}\big\|\right]^{4}\right)e^{ct}$
where $c>0$. Thus one can deduce
(45) $\mathbb{E}\left[\big\|V_{t}\big\|^{4}\right]\lesssim 1\text{ and }\mathbb{E}\left[\big\|V_{t}\big\|^{2}\right]\lesssim 1.$
Now it suffices to prove that $\mathrm{Im}T$ is in $\mathcal{C}^{1/2}[0,T]$, which is compactly embedded into $\mathcal{C}[0,T]$.
By Lemma A.2, one obtains
(46) $\big\|v_{\alpha}\left(\rho_{t}\right)-v_{\alpha}(\rho_{s})\big\|\leq c_{0}W_{2}(\rho_{t},\rho_{s}).$
For $W_{2}(\rho_{t},\rho_{s})$, it holds that
(47) $W_{2}^{2}(\rho_{t},\rho_{s})\leq\mathbb{E}\left[\big\|V_{t}-V_{s}\big\|^{2}\right].$
Further we can deduce
$V_{t}-V_{s}=\int_{s}^{t}-\lambda(V_{\tau}-u_{\tau})-\dfrac{1}{\epsilon}\nabla G(V_{\tau})\,d\tau+\sigma\int_{s}^{t}\text{diag}\left(V_{\tau}-u_{\tau}\right)\,dB_{\tau}.$
Thus
(48) $\mathbb{E}\left[\big\|V_{t}-V_{s}\big\|^{2}\right]\lesssim$ $\mathbb{E}\left[\big\|\int_{s}^{t}(V_{\tau}-u_{\tau})\,d\tau\big\|^{2}\right]+\mathbb{E}\left[\big\|\int_{s}^{t}\nabla G(V_{\tau})\,d\tau\big\|^{2}\right]$
$+\mathbb{E}\left[\big\|\int_{s}^{t}\text{diag}\left(V_{\tau}-u_{\tau}\right)\,dB_{\tau}\big\|^{2}\right].$
Now we bound from above the three terms on the right hand side respectively. For the first term, we have
(49) $\mathbb{E}\left[\big\|\int_{s}^{t}(V_{\tau}-u_{\tau})\,d\tau\big\|^{2}\right]\leq$ $\mathbb{E}\left[(\int_{s}^{t}\big\|V_{\tau}-u_{\tau}\big\|\,d\tau)^{2}\right]$
$\leq$ $|t-s|\mathbb{E}\left[\int_{s}^{t}\big\|V_{\tau}-u_{\tau}\big\|^{2}\,d\tau\right]$
$\lesssim$ $|t-s|\left(\int_{s}^{t}\mathbb{E}\left[\big\|V_{\tau}\big\|^{2}\right]\,d\tau+\int_{s}^{t}\big\|u_{\tau}\big\|^{2}\,d\tau\right)\lesssim|t-s|,$
where in the second inequality we used Cauchy’s inequality and in the last inequality, we used (45) and the fact that $u_{t}$ is continuous thus bounded in $[0,T]$.
For the second term, we have
(50) $\mathbb{E}\left[\big\|\int_{s}^{t}\nabla G(V_{\tau})\,d\tau\big\|^{2}\right]$ $\leq\mathbb{E}\left[\left(\int_{s}^{t}\big\|\nabla G(V_{\tau})\big\|\,d\tau\right)^{2}\right]$
$\leq\mathbb{E}\left[\left(\int_{s}^{t}\big\|V_{\tau}\big\|\,d\tau\right)^{2}\right]$
$\leq|t-s|\mathbb{E}\left[\int_{s}^{t}\big\|V_{\tau}\big\|^{2}\,d\tau\right]\lesssim|t-s|,$
where in the second inequality, we used Assumption 1 (4) and in the last inequality, we used (45).
For the third term in (48), we have the following estimation:
(51) $\mathbb{E}\left[\big\|\int_{s}^{t}\text{diag}\left(V_{\tau}-u_{\tau}\right)\,dB_{\tau}\big\|^{2}\right]$ $=\mathbb{E}\left[\int_{s}^{t}\big\|\text{diag}\left(V_{\tau}-u_{\tau}\right)\big\|^{2}_{F}\,d\tau\right]$
$\leq|t-s|\mathbb{E}\left[\int_{s}^{t}\big\|V_{\tau}-u_{\tau}\big\|^{4}\,d\tau\right]$
$\lesssim|t-s|(\mathbb{E}\left[\int_{s}^{t}\big\|V_{\tau}\big\|^{4}\,d\tau\right]+\int_{s}^{t}\big\|u_{\tau}\big\|^{4}\,d\tau)\lesssim|t-s|,$
where the first equality comes from Itô’s Isometry, while in the first inequality, we used Cauchy’s inequality and in the last inequality, we used (45) and the fact that $u_{t}$ is bounded.
Finally, we combine (46), (47), (48), (49), (50) and (51) to deduce
$\big\|v_{\alpha}\left(\rho_{t}\right)-v_{\alpha}(\rho_{s})\big\|\lesssim|t-s|^{1/2},$
which implies that $v_{\alpha}\left(\rho_{t}\right)\in\mathcal{C}^{0,1/2}[0,T]$. Thus, $T$ is compact.
Step 3 (Existence):
We make use of Theorem A.3. Let us take $u_{t}$ satisfying $u_{t}=\theta Tu_{t}$ for $\theta\in[0,1]$. We now try to prove $\big\|u_{t}\big\|_{\infty}\leq q$ for some finite $q>0$.
First, one has
(52) $\big\|u_{t}\big\|^{2}=\theta^{2}\big\|v_{\alpha}\left(\rho_{t}\right)\big\|^{2}\leq\theta^{2}e^{\alpha(\overline{\mathcal{E}}-\underline{\mathcal{E}})}\int\big\|v\big\|^{2}\,d\rho_{t}.$
Then, to bound $\|u_{t}\|$, we try to bound $\int\|v\|^{2}\,d\rho_{t}$. Since $\rho_{t}$ is a weak solution to the corresponding Fokker-Planck equation (8), one has
$\dfrac{d}{dt}\int\big\|v\big\|^{2}\,d\rho_{t}$ $=\int\sigma^{2}\big\|v-u_{t}\big\|^{2}-2\lambda(v-u_{t})\cdot v-\dfrac{2}{\epsilon}\nabla G(v)\cdot v\,d\rho_{t}$
$=\int(\sigma^{2}-2\lambda)\big\|v\big\|^{2}+2(\lambda-\sigma^{2})v\cdot u_{t}+d\sigma^{2}\big\|u_{t}\big\|^{2}-\dfrac{2}{\epsilon}\nabla G(v)\cdot v\,d\rho_{t}.$
Since
$\int v\cdot u_{t}\,d\rho_{t}$ $\leq\int\big\|v\big\|\cdot\big\|u_{t}\big\|\,d\rho_{t}\lesssim\int\big\|v\big\|^{2}\,d\rho_{t}+\int\big\|u_{t}\big\|^{2}\,d\rho_{t}=\int\big\|v\big\|^{2}\,d\rho_{t}+\big\|u_{t}\big\|^{2}$
and
$\big\|\nabla G(v)\big\|\lesssim\big\|v\big\|,$
one can further deduce
$\dfrac{d}{dt}\int\big\|v\big\|^{2}\,d\rho_{t}\lesssim\int\big\|v\big\|^{2}\,d\rho_{t}+\big\|u_{t}\big\|^{2}\lesssim\int\big\|v\big\|^{2}\,d\rho_{t},$
where in the last inequality, we used (52). Applying Gronwall’s inequality yields that $\int\big\|v\big\|^{2}\,d\rho_{t}$ is bounded, and from the above inequality, the bound does not depend on $u_{t}$ itself. Thus we’ have shown that $\|u_{t}\|_{\infty}$ is bounded by a uniform constant $q$. Theorem A.3 then gives the existence.
Step 4 (Uniqueness):
Suppose we are given two fixed points of $T$: $u_{t}$ and $\hat{u}_{t}$. We use $V_{t}$ and $\hat{V}_{t}$ respectively to represent the solutions of equation (44) with $u_{t}$ and $\hat{u}_{t}$ plugged in. We also assume that $V_{t}$ and $\hat{V}_{t}$ are defined in the same probability space. From the steps above, there exist constants $q>0$ and $K>0$ such that
(53) $\big\|u_{t}\big\|_{\infty},\big\|\hat{u}_{t}\big\|_{\infty}<q$
and
(54) $\sup_{t\in[0,T]}\int\big\|v\big\|^{4}\,d\rho_{t},\sup_{t\in[0,T]}\int\big\|v\big\|^{4}\,d\hat{\rho}_{t}<K,$
where $\rho_{t}$ and $\hat{\rho}_{t}$ are the distributions of $V_{t}$ and $\hat{V}_{t}$ respectively. Let us consider $Z_{t}=V_{t}-\hat{V}_{t}$. One has
$Z_{t}=$ $Z_{0}-\lambda\int_{0}^{t}Z_{\tau}\,d\tau+\lambda\int_{0}^{t}(u_{\tau}-\hat{u}_{\tau})\,d\tau-\dfrac{1}{\epsilon}\int_{0}^{t}\left(\nabla G\left(V_{\tau}\right)-\nabla G\left(\hat{V}_{\tau}\right)\right)\,d\tau$
$+\sigma\int_{0}^{t}\text{diag}\left(\left(V_{\tau}-u_{\tau}\right)-\left(\hat{V}_{\tau}-\hat{u}_{\tau}\right)\right)\,dB_{\tau}.$
Thus
(55) $\mathbb{E}\left[\big\|Z_{t}\big\|^{2}\right]$
$\lesssim$ $\mathbb{E}\left[\big\|Z_{0}\big\|^{2}\right]+\mathbb{E}\left[(\int_{0}^{t}\big\|Z_{\tau}\big\|\,d\tau)^{2}\right]+\mathbb{E}\left[(\int_{0}^{t}\big\|u_{\tau}-\hat{u}_{\tau}\big\|\,d\tau)^{2}\right]$
$\quad\quad+$ $\mathbb{E}\left[\left(\int_{0}^{t}\big\|\nabla G(V_{\tau})-\nabla G(\hat{V}_{\tau})\big\|\,d\tau\right)^{2}\right]$
$\times\mathbb{E}\left[\big\|\int_{0}^{t}\text{diag}\left(\left(V_{\tau}-u_{\tau}\right)-\left(\hat{V}_{\tau}-\hat{u}_{\tau}\right)\right)\,dB_{\tau}\big\|^{2}\right].$
For $\mathbb{E}[(\int_{0}^{t}\big\|Z_{\tau}\big\|\,d\tau)^{2}]$, we have that
(56) $\mathbb{E}\left[(\int_{0}^{t}\big\|u_{\tau}-\hat{u}_{\tau}\big\|\,d\tau)^{2}\right]$
$=\mathbb{E}\left[(\int_{0}^{t}\big\|v_{\alpha}(\rho_{\tau})-v_{\alpha}(\hat{\rho}_{\tau})\big\|\,d\tau)^{2}\right]\leq t\mathbb{E}\left[\int_{0}^{t}\big\|v_{\alpha}(\rho_{\tau})-v_{\alpha}(\hat{\rho}_{\tau})\big\|^{2}\,d\tau\right],$
where in the inequality, we used the fact that $u_{t}$ and $\hat{u}_{t}$ are fixed points. For $\mathbb{E}[(\int_{0}^{t}\big\|\nabla G(V_{\tau})-\nabla G(\hat{V}_{\tau})\big\|\,d\tau)^{2}]$, one has
(57) $\mathbb{E}\left[(\int_{0}^{t}\big\|\nabla G(V_{\tau})-\nabla G(\hat{V}_{\tau})\big\|\,d\tau)^{2}\right]$ $\lesssim\mathbb{E}\left[(\int_{0}^{t}\big\|V_{\tau}-\hat{V}_{\tau}\big\|\,d\tau)^{2}\right]$
$=\mathbb{E}\left[(\int_{0}^{t}\big\|Z_{\tau}\big\|\,d\tau)^{2}\right]\leq t\cdot\mathbb{E}\left[\int_{0}^{t}\big\|Z_{\tau}\big\|^{2}\,d\tau\right].$
Here we used the Lipschitz property of $\nabla G$. For $\mathbb{E}[\big\|\int_{0}^{t}\text{diag}\left(\left(V_{\tau}-u_{\tau}\right)-\left(\hat{V}_{\tau}-\hat{u}_{\tau}\right)\right)\,dB_{\tau}\big\|^{2}]$. Then
(58) $\mathbb{E}\left[\big\|\int_{0}^{t}\text{diag}\left(\left(V_{\tau}-u_{\tau}\right)-\left(\hat{V}_{\tau}-\hat{u}_{\tau}\right)\right)\,dB_{\tau}\big\|^{2}\right]$
$=\mathbb{E}\left[\int_{0}^{t}\big\|\text{diag}\left(\left(V_{\tau}-u_{\tau}\right)-\left(\hat{V}_{\tau}-\hat{u}_{\tau}\right)\right)\big\|^{2}_{F}\,d\tau\right]$
$\lesssim\mathbb{E}\left[\int_{0}^{t}\big\|V_{\tau}-\hat{V}_{\tau}\big\|^{2}\,d\tau\right]+\mathbb{E}\left[\int_{0}^{t}\big\|u_{\tau}-\hat{u}_{\tau}\big\|^{2}\,d\tau\right]$
$=\mathbb{E}\left[\int_{0}^{t}\big\|Z_{\tau}\big\|^{2}\,d\tau\right]+\mathbb{E}\left[\int_{0}^{t}\big\|v_{\alpha}(\rho_{\tau})-v_{\alpha}(\hat{\rho}_{\tau})\big\|^{2}\,d\tau\right],$
where in the first equality, we used Itô’s Isometry. Thus combining (55), (56), (57) and (58) yields
$\mathbb{E}\left[\big\|Z_{t}\big\|^{2}\right]\lesssim\mathbb{E}\left[\big\|Z_{0}\big\|^{2}\right]+\int_{0}^{t}\mathbb{E}\left[\big\|Z_{\tau}\big\|^{2}\right]\,d\tau+\mathbb{E}\left[\int_{0}^{t}\big\|v_{\alpha}(\rho_{\tau})-v_{\alpha}(\hat{\rho}_{\tau})\big\|^{2}\,d\tau\right].$
We further notice that by Lemma A.2,
$\big\|v_{\alpha}(\rho_{\tau})-v_{\alpha}(\hat{\rho}_{\tau})\big\|\lesssim W_{2}(\rho_{\tau},\hat{\rho}_{\tau})\leq\sqrt{\mathbb{E}\left[\big\|V_{\tau}-\hat{V}_{\tau}\big\|^{2}\right]}=\sqrt{\mathbb{E}\left[\big\|Z_{\tau}\big\|^{2}\right]}.$
So we can deduce
$\mathbb{E}\left[\big\|Z_{t}\big\|^{2}\right]$ $\lesssim\mathbb{E}\left[\big\|Z_{0}\big\|^{2}\right]+\int_{0}^{t}\mathbb{E}\left[\big\|Z_{\tau}\big\|^{2}\right]\,d\tau+\mathbb{E}\left[\int_{0}^{t}\mathbb{E}\left[\big\|Z_{\tau}\big\|^{2}\right]\,d\tau\right]$
$\lesssim\mathbb{E}\left[\big\|Z_{0}\big\|^{2}\right]+\int_{0}^{t}\mathbb{E}\left[\big\|Z_{\tau}\big\|^{2}\right]\,d\tau.$
Then applying Gronwall’s inequality with the fact that $\mathbb{E}[\big\|Z_{0}\big\|^{2}]=0$ gives the uniqueness result. ∎
A.3. Proof of Theorem 3.4
We first prove the following lemma.
Lemma A.4.
Let $\mathcal{E}$ satisfy Assumption 1 and $\rho_{0}\in\mathcal{P}_{4}(\mathbb{R}^{d})$. For any $N\geq 2$, assume that $\{(V_{t}^{i,N})_{t\in[0,T]}\}_{i=1}^{N}$ is the unique solution to the particle system (6) with $\rho_{0}^{\otimes N}$ distributed initial data $\{V_{0}^{i,N}\}_{i=1}^{N}$. Then there exists a constant $K>0$ independent of $N$ such that
$\sup_{i=1,...,N}\left\{\sup_{t\in[0,T]}\mathbb{E}\left[\big\|V_{t}^{i,N}\big\|^{2}+\big\|V_{t}^{i,N}\big\|^{4}\right]+\sup_{t\in[0,T]}\mathbb{E}\left[\big\|v_{\alpha}(\hat{\rho}_{t}^{N})\big\|^{2}+\big\|v_{\alpha}(\hat{\rho}^{N}_{t})\big\|^{4}\right]\right\}\leq K.$
Proof.
For each $i$, we have
$dV^{i,N}_{t}=-\lambda\left(V^{i,N}_{t}-v_{\alpha}(\hat{\rho}_{t})\right)\,dt-\dfrac{1}{\epsilon}\nabla G(V^{i}_{t})\,dt+\sigma\text{diag}\left(V^{i,N}_{t}-v_{\alpha}(\hat{\rho}_{t}^{i})\right)\,dB^{i}_{t},$
$V^{i}_{0}\sim\rho_{0}.$
Now we pick $p=1$ or $p=2$. Then
$\mathbb{E}\big\|V^{i,N}_{t}\big\|^{2p}\lesssim$ $\mathbb{E}\big\|V_{0}^{i,N}\big\|^{2p}+\mathbb{E}(\int_{0}^{t}\big\|V^{i,N}_{\tau}\big\|\,d\tau)^{2p}+\mathbb{E}(\int_{0}^{t}\big\|v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\big\|\,d\tau)^{2p}$
$+\mathbb{E}\big\|\int_{0}^{t}\text{diag}\left(V^{i,N}_{\tau}\right)\,dB_{\tau}^{i}\big\|^{2p}+\mathbb{E}\big\|\int_{0}^{t}\text{diag}\left(v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)\,dB_{\tau}^{i}\big\|^{2p}.$
Here, we used Assumption 1 (4). Now by Cauchy’s inequality,
$\mathbb{E}(\int_{0}^{t}\big\|V^{i,N}_{\tau}\big\|\,d\tau)^{2p}\leq t^{p}\cdot\mathbb{E}(\int_{0}^{t}\big\|V^{i,N}_{\tau}\big\|^{2}\,d\tau)^{p}$
and
$\mathbb{E}(\int_{0}^{t}\big\|v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\big\|\,d\tau)^{2p}\leq t^{p}\cdot\mathbb{E}(\int_{0}^{t}\big\|v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\big\|^{2}\,d\tau)^{p}.$
Also, by Itô Isometry,
$\mathbb{E}\big\|\int_{0}^{t}\text{diag}\left(V^{i,N}_{\tau}\right)\,dB_{\tau}^{i}\big\|^{2p}=\mathbb{E}(\int_{0}^{t}\big\|V^{i,N}_{\tau}\big\|^{2}\,d\tau)^{p}$
and
$\mathbb{E}\big\|\int_{0}^{t}\text{diag}\left(v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)\,dB_{\tau}^{i}\big\|^{2p}=\mathbb{E}(\int_{0}^{t}\big\|v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\big\|^{2}\,d\tau)^{p}.$
Thus
$\mathbb{E}\big\|V^{i,N}_{t}\big\|^{2p}\lesssim\mathbb{E}\big\|V_{0}^{i,N}\big\|^{2p}+\mathbb{E}(\int_{0}^{t}\big\|V^{i,N}_{\tau}\big\|^{2}\,d\tau)^{p}+\mathbb{E}(\int_{0}^{t}\big\|v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\big\|^{2}\,d\tau)^{p}.$
Further, by Hölder inequality,
$\mathbb{E}(\int_{0}^{t}\big\|V^{i,N}_{\tau}\big\|^{2}\,d\tau)^{p}\leq\mathbb{E}\int_{0}^{t}\big\|V^{i,N}_{\tau}\big\|^{2p}\,d\tau$
and
$\mathbb{E}(\int_{0}^{t}\big\|v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\big\|^{2}\,d\tau)^{p}\leq\mathbb{E}\int_{0}^{t}\big\|v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\big\|^{2p}\,d\tau.$
So we can deduce
$\mathbb{E}\big\|V^{i,N}_{t}\big\|^{2p}\lesssim\mathbb{E}\big\|V_{0}^{i,N}\big\|^{2p}+\mathbb{E}\int_{0}^{t}\big\|V^{i,N}_{\tau}\big\|^{2p}\,d\tau+\mathbb{E}\int_{0}^{t}\big\|v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\big\|^{2p}\,d\tau.$
Thus
(59) $\mathbb{E}\int\big\|v\big\|^{2p}\,d\hat{\rho}_{t}^{N}\lesssim\mathbb{E}\int\big\|v\big\|^{2p}\,d\hat{\rho}^{N}_{0}+\int_{0}^{t}(\mathbb{E}\int\big\|v\big\|^{2p}\,d\hat{\rho}_{\tau}^{N})\,d\tau+\int_{0}^{t}(\mathbb{E}\big\|v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\big\|^{2p})\,d\tau.$
Now by Lemma 3.3 in [12], one has
(60) $\int\big\|v\big\|^{2}\dfrac{\omega_{\alpha}(v)}{\big\|\omega_{\alpha}\big\|_{L^{1}(\tilde{\rho}_{\tau}^{N})}}\,d\hat{\rho}^{N}_{\tau}\leq b_{1}+b_{2}\int\big\|v\big\|^{2}\,d\hat{\rho}^{N}_{\tau}.$
Then we can calculate
$\big\|v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\big\|^{2p}$ $=\big\|\int v\cdot\dfrac{\omega_{\alpha}(v)}{\big\|\omega_{\alpha}\big\|_{L^{1}(\tilde{\rho}_{\tau}^{N})}}\,d\hat{\rho}^{N}_{\tau}\big\|^{2p}$
$\leq\left(\int\big\|v\big\|\cdot\dfrac{\omega_{\alpha}(v)}{\big\|\omega_{\alpha}\big\|_{L^{1}(\tilde{\rho}_{\tau}^{N})}}\,d\hat{\rho}^{N}_{\tau}\right)^{2p}$
$\leq\left(\int\big\|v\big\|^{2}\cdot\dfrac{\omega_{\alpha}(v)}{\big\|\omega_{\alpha}\big\|_{L^{1}(\tilde{\rho}_{\tau}^{N})}}\cdot\dfrac{\omega_{\alpha}(v)}{\big\|\omega_{\alpha}\big\|_{L^{1}(\tilde{\rho}_{\tau}^{N})}}\,d\hat{\rho}^{N}_{\tau}\right)^{p}$
$\leq\left(\int\big\|v\big\|^{2}\cdot\dfrac{\omega_{\alpha}(v)}{\big\|\omega_{\alpha}\big\|_{L^{1}(\tilde{\rho}_{\tau}^{N})}}\,d\hat{\rho}^{N}_{\tau}\right)^{p}$
$\leq\left(b_{1}+b_{2}\int\big\|v\big\|^{2}\,d\hat{\rho}^{N}_{\tau}\right)^{p}\lesssim 1+\int\big\|v\big\|^{2p}\,d\hat{\rho}^{N}_{\tau},$
where in the second inequality, we used Cauchy’s inequality and in the fourth inequality, we used (60) and in the last inequality, we used Hölder inequality. Combine the above inequality and (59) leads to
$\mathbb{E}\int\big\|v\big\|^{2p}\,d\hat{\rho}_{t}^{N}\lesssim\mathbb{E}\int\big\|v\big\|^{2p}\,d\hat{\rho}^{N}_{0}+\int_{0}^{t}\left(\mathbb{E}\int\big\|v\big\|^{2p}\,d\hat{\rho}_{\tau}^{N}\right)\,d\tau+1.$
By applying Gronwall’s inequality, it follows that $\mathbb{E}\int|v|^{2p}\,d\hat{\rho}_{t}^{N}$ is bounded for $t\in[0,T]$, and the bound does not depend on $N$. Also, we know that
$\big\|v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\big\|^{2p}\lesssim 1+\int\big\|v\big\|^{2p}\,d\hat{\rho}^{N}_{\tau},$
which implies that
$\mathbb{E}\big\|v_{\alpha}(\hat{\rho}^{N}_{\tau})\big\|^{2p}\lesssim 1+\mathbb{E}\int\big\|v\big\|^{2p}\,d\hat{\rho}^{N}_{t}.$
So $\mathbb{E}\big\|v_{\alpha}(\hat{\rho}^{N}_{t})\big\|^{2p}$ is bounded for $t\in[0,T]$ and the bound does not depend on $N$. ∎
As in [31], we then make the following definition.
Definition A.5.
Fix $\phi\in\mathcal{C}_{c}^{2}(\mathbb{R}^{d})$. Define functional $F_{\phi}:\mathcal{P}(\mathcal{C}[0,T];\mathbb{R}^{d})\rightarrow\mathbb{R}$:
$F_{\phi}(\mu_{t})=$ $\left\langle\phi,\mu_{t}\right\rangle-\left\langle\phi,\mu_{0}\right\rangle+\lambda\int_{0}^{t}\left\langle\left(v-v_{\alpha}(\rho_{\tau})\right)\cdot\nabla\phi(v),\mu_{\tau}\right\rangle\,d\tau$
$+\dfrac{1}{\epsilon}\int_{0}^{t}\left\langle\nabla G(v)\cdot\nabla\phi(v),\mu_{\tau}\right\rangle\,d\tau$
$-\dfrac{\sigma^{2}}{2}\int_{0}^{t}\left\langle\sum_{k=1}^{d}\left(v-v_{\alpha}(\rho_{\tau})\right)_{k}^{2}\partial_{kk}\phi(v),\mu_{\tau}\right\rangle\,d\tau.$
We can then prove the following proposition about the functional $F_{\phi}$ defined above.
Proposition A.6.
Let $\mathcal{E}$ satisfy Assumption 1 and $\rho_{0}\in\mathcal{P}_{4}(\mathbb{R}^{d})$. For any $N\geq 2$, assume that $\{(V^{i,N}_{t})\}_{i=1}^{N}$ is the unique solution to (6) with $\rho_{0}^{\otimes N}$ distributed initial data $\{V_{0}^{i,N}\}_{i=1}^{N}$. There exists a constant $C>0$ depending only on $\sigma,K,T$ and $\big\|\nabla\phi\big\|_{\infty}$ such that
$\mathbb{E}\left[|F_{\phi}(\hat{\rho}_{t}^{N})|^{2}\right]\leq\dfrac{C}{N}.$
Proof.
First we compute
$F_{\phi}(\hat{\rho}_{t}^{N})=$ $\dfrac{1}{N}\sum_{i=1}^{N}\phi(V^{i,N}_{t})-\dfrac{1}{N}\sum_{i=1}^{N}\phi(V_{0}^{i,N})$
$+\lambda\int_{0}^{t}\dfrac{1}{N}\sum_{i=1}^{N}\left(V^{i,N}_{\tau}-v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)\cdot\nabla\phi(V^{i,N}_{\tau})\,d\tau$
$+\dfrac{1}{\epsilon}\int_{0}^{t}\sum_{i=1}^{N}\nabla G(V^{i,N}_{\tau})\cdot\nabla\phi(V^{i,N}_{\tau})\,d\tau$
$-\dfrac{\sigma^{2}}{2}\int_{0}^{t}\dfrac{1}{N}\sum_{i=1}^{N}\sum_{k=1}^{d}\left(V^{i,N}_{\tau}-v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)_{kk}^{2}\partial_{kk}\phi(V^{i,N}_{\tau})\,d\tau.$
On the other hand, the Itô-Doeblin formula gives
$\phi(V^{i,N}_{t})-\phi(V_{0}^{i,N})=$ $-\lambda\int_{0}^{t}\left(V^{i,N}_{\tau}-v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)\cdot\nabla\phi(V^{i,N}_{\tau})\,d\tau$
$-\dfrac{1}{\epsilon}\int_{0}^{t}\nabla G(V^{i,N}_{\tau})\cdot\nabla\phi(V^{i,N}_{\tau})\,d\tau$
$+\sigma\int_{0}^{t}\left(\nabla\phi(V^{i,N}_{\tau})\right)^{T}\left(\text{diag}\left(V^{i,N}_{\tau}-v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)\,dB^{i}_{\tau}\right)$
$+\dfrac{\sigma^{2}}{2}\int_{0}^{t}\sum_{k=1}^{d}\left(V^{i,N}_{\tau}-v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)_{k}^{2}\partial_{kk}\phi(V^{i,N}_{\tau})\,d\tau.$
Then one gets
$F_{\phi}(\hat{\rho}_{t}^{N})=\dfrac{\sigma}{N}\int_{0}^{t}\sum_{i=1}^{N}\left(\nabla\phi(V^{i,N}_{\tau})\right)^{T}\left(\text{diag}\left(V^{i,N}_{\tau}-v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)\,dB^{i}_{\tau}\right).$
Finally, we can compute
$\mathbb{E}\left[|F_{\phi}(\hat{\rho}_{t}^{N})|^{2}\right]$ $=\dfrac{\sigma^{2}}{N^{2}}\sum_{i=1}^{N}\mathbb{E}\left[\Big|\int_{0}^{t}\sum_{i=1}^{N}\left(\nabla\phi(V^{i,N}_{\tau})\right)^{T}\text{diag}\left(V^{i,N}_{\tau}-v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)\,dB^{i}_{\tau}\Big|\right]^{2}$
$=\dfrac{\sigma^{2}}{N^{2}}\sum_{i=1}^{N}\mathbb{E}\left[\int_{0}^{t}\sum_{i=1}^{N}\big\|\left(\nabla\phi(V^{i,N}_{\tau})\right)^{T}\text{diag}\left(V^{i,N}_{\tau}-v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)\big\|_{2}^{2}\,d\tau\right]$
$\leq\dfrac{\sigma^{2}}{N^{2}}\big\|\nabla\phi\big\|_{\infty}^{2}\sum_{i=1}^{N}\int_{0}^{t}\mathbb{E}\left[\big\|V^{1,N}_{\tau}-v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\big\|_{2}^{2}\right]\,d\tau$
$\lesssim\dfrac{\sigma^{2}}{N^{2}}\big\|\nabla\phi\big\|_{\infty}^{2}\sum_{i=1}^{N}\int_{0}^{t}K\,d\tau=\dfrac{\sigma^{2}}{N^{2}}\big\|\nabla\phi\big\|_{\infty}^{2}\sum_{i=1}^{N}tK\leq T\dfrac{\sigma^{2}K}{N}\big\|\nabla\phi\big\|_{\infty},$
where in the second equality, we used Itô’s isometry and in the third inequality, we used Lemma A.4. This completes the proof. ∎
Now we recall the Aldous criteria (Section 34.3[5]), which can be used to prove the tightness of a sequence of distributions:
Lemma A.7 (The Aldous criteria).
Let $\{V^{n}\}_{n\in\mathbb{N}}$ be a sequence of random variables defined on a probability space $(\Omega,\mathcal{F},\mathbb{P})$ and valued in $\mathcal{C}([0,T];\mathbb{R}^{d})$. The sequence of probability distributions $\{\mu_{V^{n}}\}_{n\in\mathbb{N}}$ of $\{V^{n}\}_{n\in\mathbb{N}}$ is tight on $\mathcal{C}([0,T];\mathbb{R}^{d})$ if the following two conditions hold.
(Con1) For all $t\in[0,T]$, the set of distributions of $V^{n}_{t}$, denoted by $\{\mu_{V^{n}}\}_{n\in\mathbb{N}}$, is tight as a sequence of probability measures on $\mathbb{R}^{d}$.
(Con2) For all $\epsilon>0$, $\eta>0$, there exists $\delta_{0}>0$ and $n_{0}\in\mathbb{N}$ such that for all $n\geq n_{0}$ and for all discrete-valued $\sigma(V^{n}_{\tau};\tau\in[0,T])$-stopping times $\beta$ with $0\leq\beta+\delta_{0}\leq T$, it holds that
$\sup_{\delta\in[0,\delta_{0}]}\mathbb{P}\left(\big\|V^{n}_{\beta+\delta}-V^{n}_{\beta}\big\|\geq\eta\right)\leq\epsilon.$
We use the above lemma to prove the tightness of $\{\mathcal{L}(\hat{\rho}^{N})\}_{N\geq 2}$.
Theorem A.8.
Under the same assumption as in Lemma A.4, the sequence $\{\mathcal{L}(\hat{\rho}^{N})\}_{N\geq 2}$ is tight in $\mathcal{P}\left(\mathcal{P}(\mathcal{C}([0,T];\mathbb{R}^{d}))\right)$.
Proof.
It suffices to prove that $\{\mathcal{L}(V^{1,N})\}_{N\geq 2}$ is tight in $\mathcal{P}\left(\mathcal{C}([0,T];\mathbb{R}^{d})\right)$ due to Proposition 2.2(ii) in [47]. By Lemma A.7, one only needs to verify the two conditions in it. For condition 1, let us fix $\epsilon>0$. Now we consider the compact set $U_{\epsilon}=\{\big\|v\big\|^{2}\leq K/\epsilon\}$ ,where $K$ is the constant in Lemma A.4. Then by Markov’s inequality,
$\mathcal{L}(V_{t}^{1,N})(U_{\epsilon}^{c})=\mathbb{P}\left(\big\|V_{t}^{1,N}\big\|>\dfrac{\epsilon}{K}\right)\leq\dfrac{\epsilon\mathbb{E}\left[\big\|V^{1,N}_{t}\big\|^{2}\right]}{K}\leq\epsilon$
for any $N\geq 2$, where in the last inequality we used Lemma A.4. Thus condition 1 is verified.
For condition 2, we fix $\epsilon>0$ and $\eta>0$. Notice that
(61) $V^{1,N}_{\beta+\delta}-V^{1,N}_{\beta}=$ $-\lambda\int_{\beta}^{\beta+\delta}\left(V_{\tau}^{1,N}-v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)\,d\tau+\sigma\int_{\beta}^{\beta+\delta}\text{diag}\left(V^{1,N}_{\tau}-v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)\,dB_{\tau}^{1}$
$-\dfrac{1}{\epsilon}\int_{\beta}^{\beta+\delta}\nabla G(V^{1,N}_{\tau})\,d\tau.$
Following the same steps in the proof of Lemma 2.1 in [31],
(62) $\mathbb{E}\left[\big\|\lambda\int_{\beta}^{\beta+\delta}\left(V^{1,N}_{\tau}-v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)\,d\tau\big\|^{2}\right]\leq 2TK\lambda^{2}\delta$
and
(63) $\mathbb{E}\left[\big\|\sigma\int_{\beta}^{\beta+\delta}\text{diag}\left(V^{1,N}-v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)\,dB_{\tau}^{1}\big\|^{2}\right]\leq\sigma^{2}\sqrt{8\delta TK}.$
Also, we can compute
$\big\|\dfrac{1}{\epsilon}\int_{\beta}^{\beta+\delta}\nabla G(V^{1,N}_{\tau})\,d\tau\big\|^{2}$ $\leq\dfrac{1}{\epsilon^{2}}\int_{\beta}^{\beta+\delta}\big\|\nabla G(V^{1,N}_{\tau})\big\|\,d\tau$
$\lesssim\left(\int_{\beta}^{\beta+\delta}\big\|V^{i,N}_{\tau}\big\|\,d\tau\right)^{2}\leq\delta\int_{\beta}^{\beta+\delta}\big\|V^{i,N}_{\tau}\big\|^{2}\,d\tau,$
where in the second inequality we used Assumption 1 (4). Thus
$\mathbb{E}\left[\big\|\dfrac{1}{\epsilon}\int_{\beta}^{\beta+\delta}\nabla G(V^{1,N}_{\tau})\,d\tau\big\|^{2}\right]$ $\lesssim\delta\mathbb{E}\left[\int_{\beta}^{\beta+\delta}\big\|V^{i,N}_{\tau}\big\|^{2}\,d\tau\right]$
$=\delta\int_{\beta}^{\beta+\delta}\mathbb{E}\big\|V^{i,N}_{\tau}\big\|^{2}\,d\tau\lesssim\delta\int_{\beta}^{\beta+\delta}\,d\tau=\delta^{2}$
where we used Lemma A.4. Combining the above inequality and (61), (62) and (63), we can conclude
$\mathbb{E}\left[\big\|V_{\beta}^{1,N}-V_{\beta+\delta}^{1,N}\big\|^{2}\right]\lesssim O(\sqrt{\delta}).$
Then one can deduce
$\mathbb{P}(\big\|V_{\beta}^{1,N}-V_{\beta+\delta}^{1,N}\big\|>\eta)\leq\dfrac{\mathbb{E}\left[\big\|V_{\beta}^{1,N}-V_{\beta+\delta}^{1,N}\big\|^{2}\right]}{\eta}\lesssim\dfrac{O(\sqrt{\delta})}{\eta}.$
Choose $\delta_{0}$ small enough finishes the proof. ∎
By Shorokhod’s lemma, for every convergent subsequence of $\{\hat{\rho}_{t}^{N}\}_{N\in\mathbb{N}}$, which is denoted by the sequence itself for simplicity and has $\rho_{t}$ as limit, one can find a probability space space $(\Omega,\mathcal{F},\mathbb{P})$ on which $\hat{\rho}_{t}^{N}$ converges to $\rho_{t}$ as random variables valued in $\mathcal{P}(\mathcal{C}[0,T];\mathbb{R}^{d})$. We use $V_{N}$ to denote the corresponding random variable of $\hat{\rho}_{t}^{N}$ and $V$ to denote the corresponding random variable of $\rho_{t}$. Moreover, by the dominated convergence theorem, one has
(64) $\left\langle\phi,\hat{\rho}_{t}^{N}-\rho_{t}\right\rangle\rightarrow 0$
almost surely for fixed $t\in[0,T]$ and $\phi\in\mathcal{C}_{b}(\mathbb{R}^{d})$.
After all these preparations, we now prove Theorem 3.4.
Proof of Theorem 3.4.
We first show that every convergent sequence converges to a solution of (8). Now suppose we have a convergent subsequence of $\{\hat{\rho}_{t}^{N}\}_{N\in\mathbb{N}}$, which is denoted by the sequence itself for simplicity and has $\rho_{t}$ as limit. Also, we use $V_{N}$ and $V$ to denote the corresponding random variables generated by Shorokhod’s lemma as mentioned above. We verify that $\rho_{t}$ is a solution to the Fokker-Planck equation (8).
For continuity, we have that for any $\phi\in\mathcal{C}_{c}^{2}(\mathbb{R}^{d})$ and $t_{n}\rightarrow t$:
$\left\langle\phi,\rho_{t_{n}}\right\rangle=\int\phi\left(V(t_{n})\right)\,d\mathbb{P}\rightarrow\int\phi\left(V(t)\right)\,d\mathbb{P}=\left\langle\phi,\rho_{t}\right\rangle.$
To prove $\rho_{t}$ satisfies the Fokker-Planck equation (8), we first prove the following four limits:
  1. (1)
    $\mathbb{E}\left[\left(\left\langle\phi,\hat{\rho}_{t}^{N}\right\rangle-\left\langle\phi,\hat{\rho}_{0}\right\rangle\right)-\left(\left\langle\phi,\rho_{t}\right\rangle-\left\langle\phi,\rho_{0}\right\rangle\right)\right]$ converges to 0 as $N\rightarrow\infty$.
  2. (2)
    $\mathbb{E}\left[\int_{0}^{t}\left\langle\left(v-v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)\cdot\nabla\phi(v),\hat{\rho}_{\tau}^{N}\right\rangle\,d\tau-\int_{0}^{t}\left\langle\left(v-v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)\cdot\nabla\phi(v),\rho_{\tau}\right\rangle\,d\tau\right]$ converges to 0 as $N\rightarrow\infty$.
  3. (3)
    The quantity
    $\mathbb{E}\Bigg[$ $\int_{0}^{t}\left\langle\sum_{k=1}^{d}\left(v-v_{\alpha}\big(\hat{\rho}^{N}_{\tau}\big)\right)_{k}^{2}\partial_{kk}\phi(v),\hat{\rho}_{\tau}^{N}\right\rangle\,d\tau$
    $-\int_{0}^{t}\left\langle\sum_{k=1}^{d}\left(v-v_{\alpha}(\rho_{\tau})\right)_{k}^{2}\partial_{kk}\phi(v),\rho_{\tau}\right\rangle\,d\tau\Bigg]$
    converges to 0 as $N\rightarrow\infty$.
  4. (4)
    $\mathbb{E}\left[\int_{0}^{t}\left\langle\nabla G(v)\cdot\nabla\phi(v),\hat{\rho}_{\tau}^{N}\right\rangle\,d\tau-\int_{0}^{t}\left\langle\nabla G(v)\cdot\nabla\phi(v),\rho_{\tau}\right\rangle\,d\tau\right]$ converges to 0 as $N\rightarrow\infty$.
The first three limits can be proved using the same methods as in Theorem 3.3 in [31] and the last one is a direct result of (64). Combining the above four limits gives
$\mathbb{E}\left[F_{\phi}(\rho_{t})-F_{\phi}(\hat{\rho}_{t}^{N})\right]=0.$
Then we can deduce
$\Big|\mathbb{E}\left[F_{\phi}(\rho_{t})\right]\Big|\leq\lim_{N\rightarrow\infty}\Big|\mathbb{E}\left[F_{\phi}(\rho_{t})-F_{\phi}(\hat{\rho}_{t}^{N})\right]\Big|+\Big|\mathbb{E}\left[F_{\phi}(\hat{\rho}_{t}^{N})\right]\Big|\leq 0+\lim_{N\rightarrow\infty}\sqrt{\dfrac{C}{N}}=0,$
where in the last inequality, we used Proposition A.6. Thus $F_{\phi}(\rho_{t})=0$ almost surely, which implies that $\rho_{t}$ is a solution to the corresponding Fokker-Planck equation (8).
Then we utilize Lemma A.10 to establish that every convergent subsequence converges to the same limit: the unique solution to (8). Combining with Theorem A.8, we deduce that $\{\hat{\rho}_{t}^{N}\}_{N\in\mathbb{N}}$ converges and the limit is exactly the solution to (8). ∎
A.4. Some auxiliary results used in the proof of Theorem 3.4
Theorem A.9.
$\forall T>0$, let $b_{t}\in\mathcal{C}\left([0,T];\mathbb{R}^{d}\right)$ and $\rho_{0}\in\mathcal{P}_{2}(\mathbb{R}^{d})$. The following linear PDE
(65) $\partial_{t}\rho_{t}=\lambda\nabla\cdot\left(\left((v-b_{t})+\dfrac{1}{\epsilon}\nabla G(v)\right)\rho_{t}\right)+\dfrac{\sigma^{2}}{2}\sum_{k=1}^{d}\partial_{x_{k}x_{k}}\left((v-b_{t})_{k}^{2}\rho_{t}\right)$
has a unique weak solution $\rho_{t}\in\mathcal{C}([0,T];\mathcal{P}_{2}(\mathbb{R}^{d}))$.
Proof.
We can obtain a solution to (65) as the law of the solution to the associated linear SDE to (65). Thus we have the existence result. For uniqueness, let us fix $t_{0}\in[0,T]$ and $\psi\in\mathcal{C}_{c}^{\infty}(\mathbb{R}^{d})$. We then can solve the following backward PDE
$\partial_{t}h_{t}=\left(\lambda(v-b_{t})+\dfrac{1}{\epsilon}\nabla G(v)\right)\cdot\nabla h_{t}-\dfrac{\sigma^{2}}{2}\sum_{k=1}^{d}(v-b_{t})_{k}^{2}\partial_{x_{k}x_{k}}h_{t},$
$(t,v)\in[0,t_{0}]\times\mathbb{R}^{d};h_{t_{0}}=\psi.$
It has a classical solution:
$h_{t}=\mathbb{E}\left[\psi(V_{t_{0}}^{t,v})\right],t\in[0,t_{0}],$
where $(V_{\tau}^{t,x})_{0\leq t\leq s\leq t_{0}}$ is the strong solution to
$dV_{\tau}^{t,v}=-\left(\lambda(V_{\tau}^{t,v}-b_{\tau})+\dfrac{1}{\epsilon}\nabla G(V_{\tau}^{t,v})\right)\,d\tau+\sigma\text{diag}\left(V_{\tau}^{t,v}-b_{\tau}\right)\,dB_{\tau},V_{t}^{t,v}=v.$
Suppose $\rho^{1}$ and $\rho^{2}$ are two weak solutions to (65). Consider $\delta\rho=\rho^{1}-\rho^{2}$. Then
$\left\langle h_{t_{0}},\delta\rho_{t_{0}}\right\rangle$ $=$ $\int_{0}^{t_{0}}\left\langle\partial_{\tau}h_{\tau},\delta\rho_{\tau}\right\rangle\,d\tau-\lambda\int_{0}^{t_{0}}\left\langle(v-b_{\tau})\nabla h_{\tau},\delta\rho_{\tau}\right\rangle\,d\tau$
$-\dfrac{1}{\epsilon}\int_{0}^{t_{0}}\left\langle\nabla G\cdot\nabla h_{\tau},\delta\rho_{\tau}\right\rangle\,d\tau+\dfrac{\sigma^{2}}{2}\int_{0}^{t_{0}}\left\langle\sum_{k=1}^{d}(v-b_{\tau})_{k}^{2}\partial_{kk}h_{\tau},\delta\rho_{\tau}\right\rangle\,d\tau$
$=$ $\int_{0}^{t_{0}}\left\langle\partial_{\tau}h_{\tau},\delta\rho_{\tau}\right\rangle\,d\tau+\int_{0}^{t_{0}}\left\langle-\partial_{\tau}h_{\tau},\delta\rho_{\tau}\right\rangle\,d\tau=0.$
This implies that $\int\psi\delta\rho_{t_{0}}=0$ for any chosen $\psi\in\mathcal{C}_{c}^{\infty}(\mathbb{R}^{d})$ and $t_{0}\in[0,T]$. Thus $\delta\rho_{t}=0.$ This proves the uniqueness. ∎
Lemma A.10.
Assume that $\rho^{1},\rho^{2}\in\mathcal{C}\left([0,T];\mathcal{P}_{2}(\mathbb{R}^{d})\right)$ are two weak solutions to PDE (8) in the sense of Definition 3.2 with the same initial data $\rho_{0}$. Then it holds that
$\sup_{t\in[0,T]}W_{2}\left(\rho^{1}_{t},\rho^{2}_{t}\right)=0,$
where $W_{2}$ is the 2-Wasserstein distance.
Proof.
Given $\rho^{1}$ and $\rho^{2}$, we first solve the following two linear SDEs
$d\tilde{V}^{i}_{t}=-\lambda\left(\tilde{V}^{i}_{t}-v_{\alpha}(\rho_{t}^{i})\right)\,dt-\dfrac{1}{\epsilon}\nabla G\,dt+\sigma\text{diag}\left(\tilde{V}^{i}_{t}-v_{\alpha(\rho_{t}^{i})}\right)\,dB_{t},$
$\hat{V}^{i}_{0}\sim\rho_{0}$
for $i=1,2$. We use $\tilde{\rho}_{t}^{i}$ to denote the law of $\tilde{V}^{i}_{t}$ for $i=1,2$. Thus $\tilde{\rho}_{t}^{i}$ solves
$\begin{gathered}\partial_{t}\tilde{\rho}_{t}^{i}=\lambda\text{div}\left((v-v_{\alpha}(\rho_{t}^{i})+\dfrac{1}{\epsilon}\nabla G)\tilde{\rho}_{t}^{i}\right)+\dfrac{\sigma^{2}}{2}\sum_{k=1}^{d}\partial_{x_{k}x_{k}}\left(\big\|v-v_{\alpha}(\rho_{t}^{i})\big\|^{2}\tilde{\rho}_{t}^{i}\right),\\ \tilde{\rho}^{i}_{0}=\rho_{0}\end{gathered}$
in the weak sense for $i=1,2$. Moreover, $\rho^{i}$ solves the above PDE since we assumed that $\rho^{i}$ solves (8). But from Theorem A.10, the solution to the above PDE is unique for $i=1,2$. This implies that $\tilde{\rho}^{i}_{t}=\rho^{i}_{t}$ for $i=1,2$. As a result, $\tilde{V}^{1}_{t}$ and $\tilde{V}^{2}_{t}$ both solve (7). By Theorem 3.3, it holds that
$\sup_{t\in[0,T]}\mathbb{E}\left[|\tilde{V}^{1}_{t}-\tilde{V}^{2}_{t}|^{2}\right]=0.$
Then one has
$\sup_{t\in[0,T]}W_{2}^{2}\left(\rho^{1},\rho^{2}\right)=\sup_{t\in[0,T]}W_{2}^{2}\left(\tilde{\rho}^{1},\tilde{\rho}^{2}\right)\leq\sup_{t\in[0,T]}\mathbb{E}\left[|\tilde{V}^{1}_{t}-\tilde{V}^{2}_{t}|^{2}\right]=0$
This completes the proof. ∎
Appendix B Proof of Lemma 4.2
Proof.
We verify Assumption 2 (B) one by one.
Verification of Assumption 2 (B1)
We first prove when $K$ is convex, the distance function $g$ is convex. Given $\lambda_{1},\lambda_{2}>0$ with $\lambda_{1}+\lambda_{2}=1$ and $v_{1},v_{2}\in\mathbb{R}^{d}$. By the definition of distance function, for any $\epsilon>0$, there exist $u_{1}$ and $u_{2}$ in $K$ such that,
$\|v_{1}-u_{1}\|_{2}\leq g(v_{1})+\epsilon,$
$\|v_{2}-u_{2}\|_{2}\leq g(v_{2})+\epsilon.$
Thus one has
$g(\lambda_{1}v_{1}+\lambda_{2}v_{2})$ $\leq\|\lambda_{1}v_{1}+\lambda_{2}v_{2}-\lambda_{1}u_{1}-\lambda_{2}u_{2}\|_{2}$
$\leq\lambda_{1}\|v_{1}-u_{1}\|_{2}+\lambda_{2}\|v_{2}-u_{2}\|_{2}$
$\leq\lambda_{1}g(v_{1})+\lambda_{2}g(v_{2})+\epsilon.$
The above inequality holds for all $\epsilon>0$. This gives the convexity of $g$. Also, one knows that $x^{2}$ is a convex increasing function when $x\geq 0$. As a result, the composition $G=g^{2}$ is convex, thereby satisfying Assumption 2 (B1).
Verification of Assumption 2 (B2)
Since we are dealing with a compact convex set $K\subset\mathbb{R}^{d}$, we can conveniently work with the Moreau envelope. Let $P_{K}$ denote the metric projection onto $K$. Then
$G(v)=\operatorname{dist}(v,K)^{2}=\|v-P_{K}(v)\|^{2},\qquad v\in\mathbb{R}^{d}.$
Let $\iota_{K}$ be the indicator function of $K$, that is,
$\iota_{K}(v)=\begin{cases}0,&v\in K,\\ +\infty,&v\notin K.\end{cases}$
The Moreau envelope of $\iota_{K}$ is defined by
$M_{\iota_{K}}(u)=\inf_{v\in\mathbb{R}^{d}}\biggl\{\iota_{K}(v)+\frac{1}{2}\,\|v-u\|^{2}\biggr\}.$
Since $\iota_{K}(v)=0$ for $v\in K$ and $\infty$ otherwise, we obtain
$M_{\iota_{K}}(u)=\inf_{v\in K}\frac{1}{2}\,\|v-u\|^{2}=\frac{1}{2}\,\|u-P_{K}(u)\|^{2}=\frac{1}{2}G(u),$
so $M_{\iota_{K}}=G/2$.
We now use [6, Proposition 12.30], which yields
$\nabla M_{\iota_{K}}(u)=u-\mathrm{prox}_{\iota_{K}}(u),$
where $\mathrm{prox}_{\iota_{K}}$ is the proximal operator of $\iota_{K}$. In the present case one has $\mathrm{prox}_{\iota_{K}}=P_{K}$, and therefore
$\nabla G(v)=2\nabla M_{\iota_{K}}(v)=2\bigl(v-P_{K}(v)\bigr),$
so that
$\|\nabla G(v)\|^{2}=4\|\nabla M_{\iota_{K}}(v)\|^{2}=4\|v-P_{K}(v)\|^{2}=4G(v),$
for all $v\in\mathbb{R}^{d}$. Hence the inequality in Assumption 2 (B2) is satisfied with equality.
It remains to prove that $G\in\mathcal{C}^{2}_{*}(\mathbb{R}^{d})$. In other words, we need to show that $|\partial_{k}G(x)|\leq C(1+|x_{k}|)$ and $\sup_{x\in\mathbb{R}^{d}}|\partial_{kk}G(x)|<\infty$ for all $k=1,\dots,d$.
The first bound follows immediately from the expression $\nabla G(v)=2\bigl(v-P_{K}(v)\bigr)$:
$|\partial_{k}G(v)|=2\,|v_{k}-P_{K}(v)_{k}|\leq 2\bigl(|v_{k}|+\sup_{u\in K}|u_{k}|\bigr)\leq C(1+|v_{k}|),$
where $\sup_{u\in K}|u_{k}|<\infty$ because $K$ is compact, and $C>0$ is a constant independent of $v$ and of the dimension.
For the second bound, we recall that the gradient of a Moreau envelope is always globally Lipschitz; in particular, $\nabla M_{\iota_{K}}$ is 11–Lipschitz, hence $\nabla G=2\nabla M_{\iota_{K}}$ is globally Lipschitz. By Rademacher’s theorem, $G$ is twice differentiable almost everywhere, and the Hessian $\nabla^{2}G(x)$ exists for almost every $x$ with operator norm bounded by the Lipschitz constant of $\nabla G$. Thus $\|\nabla^{2}G(x)\|_{\mathrm{op}}$ is essentially bounded, and the diagonal entries $\partial_{kk}G(x)$ (which are nonnegative due to the convexity of $G$) are uniformly bounded as well. This yields
$\sup_{x\in\mathbb{R}^{d}}|\partial_{kk}G(x)|<\infty,\qquad k=1,\dots,d.$
Hence $G\in\mathcal{C}^{2}_{*}(\mathbb{R}^{d})$, and Assumption 2 (B2) is verified. Moreover, all the constants involved can be chosen independently of the ambient dimension.
Verification of Assumption 2 (B3)
We first recall the relevant notions from convex geometry.
Definition B.1 (Intrinsic volumes, [34, Definition 1.1]).
Let $S\subset\mathbb{R}^{d}$ be a nonempty compact convex set. For each index $j=0,1,\dots,d$, let $P_{j}\in\mathbb{R}^{d\times d}$ be the orthogonal projector onto a fixed jj–dimensional subspace of $\mathbb{R}^{d}$, and let $Q\in\mathbb{R}^{d\times d}$ be a random rotation matrix distributed according to the Haar measure on the special orthogonal group $\mathrm{SO}(d)$. The intrinsic volumes of $S$ are defined by
$V_{j}(S):=\binom{d}{j}\frac{\kappa_{d}}{\kappa_{j}\kappa_{d-j}}\,\mathbb{E}_{Q}\bigl[\operatorname{Vol}_{j}(P_{j}QS)\bigr],\qquad j=0,\dots,d,$
where $\kappa_{d}$ denotes the volume of the dd–dimensional Euclidean unit ball and $\operatorname{Vol}_{j}(\cdot)$ is the jj–dimensional Lebesgue measure computed relative to the affine hull. In particular, $V_{d}(S)$ coincides with the dd–dimensional Lebesgue measure of $S$.
Definition B.2 (Parallel sets).
Let $K\subset\mathbb{R}^{d}$ be a nonempty compact convex set and let $\operatorname{dist}(\cdot,K)$ denote the Euclidean distance to $K$. For $r\geq 0$ the parallel set of $K$ at distance $r$ is defined by
$K_{\leq r}:=\{v\in\mathbb{R}^{d}:\operatorname{dist}(v,K)\leq r\}.$
We will use the following standard results.
Proposition B.3 (Steiner formula, [34, Fact 2.1]).
Let $K\subset\mathbb{R}^{d}$ be a nonempty compact convex set. For every $r\geq 0$ one has
$V_{d}(K_{\leq r})=\sum_{j=0}^{d}\kappa_{d-j}V_{j}(K)\,r^{d-j}.$
Proposition B.4 ([42, Corollary 2.5]).
Let $K\subset\mathbb{R}^{d}$ be a nonempty compact convex set. For all $r>0$, except possibly for a countable set of values of $r$, one has
$\frac{d}{dr}V_{d}(K_{\leq r})=\mathcal{H}^{d-1}(\partial K_{\leq r}),$
where $\mathcal{H}^{d-1}(\cdot)$ denotes the (d1)(d-1)–dimensional Hausdorff measure.
We now prove Assumption 2 (B3) holds. Recall that $G=g^{2}$ with $g(v)=\operatorname{dist}(v,K)$, and from the verification of (B2) we have
$\|\nabla G(v)\|=\sqrt{4G(v)}=2g(v),\qquad v\in\mathbb{R}^{d}.$
At points where $g$ is differentiable and $g(v)>0$,
$\nabla G(v)=2g(v)\,\nabla g(v),$
so
$2g(v)=\|\nabla G(v)\|=\|2g(v)\,\nabla g(v)\|=2g(v)\,\|\nabla g(v)\|,$
and hence $\|\nabla g(v)\|=1$ at all such points.
By the coarea formula applied to the Lipschitz function $g$ we obtain
$\int_{\{G\in(0,1)\}}\frac{1}{\|\nabla G(v)\|}\,dv$ $=\int_{\{G\in(0,1)\}}\frac{1}{2g(v)}\,\|\nabla g(v)\|\,dv$
$=\int_{0}^{1}\int_{\{g(v)=t\}}\frac{1}{2g(v)}\,d\mathcal{H}^{d-1}(v)\,dt$
$=\int_{0}^{1}\frac{1}{2t}\,\mathcal{H}^{d-1}\bigl(\{v\in\mathbb{R}^{d}:g(v)=t\}\bigr)\,dt$
$=\int_{0}^{1}\frac{1}{2t}\,\mathcal{H}^{d-1}(\partial K_{\leq t})\,dt,$
where in the last step we used $\{v:g(v)=t\}=\partial K_{\leq t}$.
From Propositions B.3 and B.4 we have, for all $t>0$ outside a countable set,
$\mathcal{H}^{d-1}(\partial K_{\leq t})$ $=\frac{d}{dt}V_{d}(K_{\leq t})$
$=\sum_{j=0}^{d-1}(d-j)\,\kappa_{d-j}\,V_{j}(K)\,t^{d-j-1}.$
Therefore,
$\int_{\{G\in(0,1)\}}\frac{1}{\|\nabla G(v)\|}\,dv=\frac{1}{2}\sum_{j=0}^{d-1}(d-j)\,\kappa_{d-j}\,V_{j}(K)\int_{0}^{1}t^{d-j-2}\,dt.$
For $j\leq d-2$ the integral $\int_{0}^{1}t^{d-j-2}\,dt$ is finite, whereas for j=d1j=d-1 the integrand behaves like $t^{-1}$ near 0 and the integral diverges. Hence finiteness of the above expression requires
$V_{d-1}(K)=0.$
A sufficient condition for this is that $K$ is contained in an affine subspace of dimension less than or equal to d2d-2, in which case the (d1)(d-1)–dimensional intrinsic volume $V_{d-1}(K)$ vanishes. Under this condition, Assumption 2 (B3) holds.
Appendix C Proof of Lemma 4.4
Proof.
We first prove the existence of $\tau_{2}$. To begin with, one deduces
$|\underline{\mathcal{E}_{u}}|=|\mathcal{E}(v_{u})|=|\mathcal{E}(v_{u})-0|=|\mathcal{E}(v_{u})-\mathcal{E}(v^{*})|\leq C\big\|v_{u}-v^{*}\big\|_{\infty}^{\beta}\leq C\tau_{1}^{\beta}(u),$
where the first inequality comes from Assumption 2 (A2) and the second inequality comes from Assumption 2 (C1). Then by taking $\tau_{2}(x)$ to be $\tau_{1}^{\beta}(x)$ will finish the proof of the existence of $\tau_{2}$.
For the existence of $\tau_{3}$, we can first pick $v_{1}\in B^{\infty}(v_{u},r)\cap\{G=u\}$, $v_{2}\in B^{\infty}(v^{*},r)$ and then do the following calculation:
$|\mathcal{E}(v_{1})-\mathcal{E}(v_{2})|$ $\leq C\big\|v_{1}-v_{2}\big\|_{\infty}^{\beta}$
$\lesssim(\big\|v_{1}-v_{u}\big\|_{\infty}^{\beta}+\big\|v_{u}-v^{*}\big\|_{\infty}^{\beta}+\big\|v^{*}-v_{2}\big\|_{\infty}^{\beta})$
$\leq(r^{\beta}+\tau_{1}(u)^{\beta}+r^{\beta})$
$\leq\left(\max\{u,r\}^{\beta}+\tau_{1}(\max\{u,r\})^{\beta}+\max\{u,r\}^{\beta}\right),$
where in the first inequality, we used Assumption 2 (A2) and in the third inequality, we used Assumption 2 (C1). Then one has
$\sup_{v_{1}\in B(v_{u},r)\cap\{G=u\},\atop v_{2}\in B(v^{*},r)}|\mathcal{E}(v_{1})-\mathcal{E}(v_{2})|\lesssim\left(\max\{u,r\}^{\beta}+\tau_{1}(\max\{u,r\})^{\beta}+\max\{u,r\}^{\beta}\right).$
So
$|\mathcal{E}_{r}^{u}-\mathcal{E}_{r}|\lesssim\left(\max\{u,r\}^{\beta}+\tau_{1}(\max\{u,r\})^{\beta}+\max\{u,r\}^{\beta}\right).$
Therefore, selecting $\tau_{3}(x)$ as a scalar multiple of $2x^{\beta}+\tau_{1}^{\beta}(x)$ will suffice. One can apply the same method to prove the existence of $\tau_{4}(x)$. ∎
Appendix D Explanations for expanding the test function space
We follow the same argument as in [23]. To start with, for any $\phi\in\mathcal{C}_{*}^{2}(\mathbb{R}^{d})$, one apply Itǒ’s formula to $\bar{V}_{t}$ to get
$d\phi(\bar{V}_{t})=$ $\nabla\phi\left(\bar{V}_{t}\right)\cdot\left(\left(-\lambda\left(\bar{V}_{t}-v_{\alpha}\big(\rho_{t}\big)\right)-\dfrac{1}{\epsilon}\nabla G\left(\bar{V}_{t}\right)\right)\,dt\right)$
$+\dfrac{1}{2}\sigma^{2}\sum_{k=1}^{d}\partial_{kk}\phi\left(\bar{V}_{t}\right)\left(\bar{V}_{t}-v_{\alpha}\big(\rho_{t}\big)\right)_{k}^{2}\,dt+\sigma\nabla\phi\left(\bar{V}_{t}\right)^{T}\text{diag}\left(\bar{V}_{t}-v_{\alpha}\big(\rho_{t}\big)\right)\,dB_{t}.$
Note that $\mathbb{E}\int_{0}^{t}\sigma\nabla\phi\left(\bar{V}_{t}\right)^{T}\text{diag}\left(\bar{V}_{t}-v_{\alpha}\big(\rho_{t}\big)\right)\,dB_{t}=0$ by applying Theorem 3.2.1 (iii) in [39] due to the facts that $\phi\in\mathcal{C}_{*}^{2}\left(\mathbb{R}^{d}\right)$ and $\rho_{t}\in\mathcal{C}\left([0,T],\mathcal{P}_{4}\left(\mathbb{R}^{d}\right)\right)$. Taking the expectation and applying Fubini’s theorem gives
$\dfrac{d}{dt}\mathbb{E}\phi\left(\bar{V}_{t}\right)=$ $-\lambda\mathbb{E}\nabla\phi\left(\bar{V}_{t}\right)\cdot\left(-\lambda\left(\bar{V}_{t}-v_{\alpha}\big(\rho_{t}\big)\right)-\dfrac{1}{\epsilon}\nabla G\left(\bar{V}_{t}\right)\right)$
$+\dfrac{1}{2}\sigma^{2}\mathbb{E}\sum_{k=1}^{d}\partial_{kk}\phi\left(\bar{V}_{t}\right)\left(\bar{V}_{t}-v_{\alpha}\big(\rho_{t}\big)\right)_{k}^{2},$
which is exactly Definition 3.2 (ii) with $\phi$ being a function in $\mathcal{C}_{*}^{2}\left(\mathbb{R}^{d}\right)$.
Appendix E Proof of Lemma 4.3
Proof.
Substituting $\phi(v)=\dfrac{1}{2}\|v-v^{*}\|^{2}$ into Definition 3.2 gives
(66) $\dfrac{d}{dt}\mathcal{V}\left(\rho_{t}\right)=$ $-\lambda\int\Bigl\langle v-v_{\alpha}\left(\rho_{t}\right),v-v^{*}\Bigr\rangle\,d\rho_{t}$
$-\dfrac{1}{\epsilon}\int\Bigl\langle\nabla G,v-v^{*}\Bigr\rangle\,d\rho_{t}+\dfrac{\sigma^{2}}{2}\int\|v-v(\rho_{t})\|^{2}\,d\rho_{t}.$
Notice that
$-\lambda\int\Bigl\langle v-v_{\alpha}\left(\rho_{t}\right),v-v^{*}\Bigr\rangle\,d\rho_{t}(v)$
$=-\lambda\int\Bigl\langle v-v^{*},v-v^{*}\Bigr\rangle\,d\rho_{t}(v)+\lambda\int\Bigl\langle v-v^{*},v_{\alpha}\big(\rho_{t}\big)-v^{*}\Bigr\rangle\,d\rho_{t}(v)$
$=-2\lambda\mathcal{V}\left(\rho_{t}\right)+\lambda\Bigl\langle\int(v-v^{*})\,d\rho_{t}(v),v_{\alpha}\left(\rho_{t}\right)-v^{*}\Bigr\rangle.$
Then one can deduce
(67) $-\lambda\int\Bigl\langle v-v_{\alpha}\left(\rho_{t}\right),v-v^{*}\Bigr\rangle\,d\rho_{t}(v)$
$\leq-2\lambda\mathcal{V}\left(\rho_{t}\right)+\lambda\|\int(v-v^{*})\,d\rho_{t}(v)\|_{2}\cdot\|v_{\alpha}\left(\rho_{t}\right)-v^{*}\|_{2}$
$\leq-2\lambda\mathcal{V}\left(\rho_{t}\right)+\lambda\int\|(v-v^{*})\|_{2}\,d\rho_{t}(v)\cdot\|v_{\alpha}\left(\rho_{t}\right)-v^{*}\|_{2}$
$\leq-2\lambda\mathcal{V}\left(\rho_{t}\right)+\lambda\sqrt{2\mathcal{V}\left(\rho_{t}\right)}\cdot\|v_{\alpha}\left(\rho_{t}\right)-v^{*}\|_{2},$
where the first and third inequalities come from Cauchy’s inequality and the second inequality is a consequence of Minkowski’s inequality.
For the last term on the right-hand side of (66), we can do the following estimate,
(68) $\dfrac{\sigma^{2}}{2}\int\|v-v(\rho_{t})\|_{2}^{2}\,d\rho_{t}(v)$
$=$ $\dfrac{\sigma^{2}}{2}\left(\int\|v-v^{*}\|_{2}^{2}\,d\rho_{t}(v)-2\Bigl\langle\int(v-v^{*})\,d\rho_{t}(v),v_{\alpha}\left(\rho_{t}\right)-v^{*}\Bigr\rangle+\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{2}^{2}\right)$
$\leq$ $\sigma^{2}\left(\mathcal{V}\big(\rho_{t}\big)+\int\|v-v^{*}\|_{2}\,d\rho_{t}(v)\cdot\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{2}+\dfrac{1}{2}\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{2}^{2}\right)$
$\leq$ $\sigma^{2}\left(\mathcal{V}\big(\rho_{t}\big)+\sqrt{2\mathcal{V}\big(\rho_{t}\big)}\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{2}+\dfrac{1}{2}\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{2}^{2}\right),$
where in the first inequality, we use Cauchy’s inequality and Minkowski’s inequality and in the second inequality, we use Cauchy’s inequality again. Plugging (67) and (68) back into (66) finishes the proof. ∎
Appendix F Lemmas used in Laplace’s principle
In this section, we present the two lemmas (Lemma F.1 and Lemma F.2) used in the proof of Lemma 4.5.
F.1. Lemma F.1 and proof
Lemma F.1.
Fix $r\in(0,R_{0})$ small enough. $\forall q>0$ with $q+\mathcal{E}^{0}_{r}<\mathcal{E}_{\infty}$,
$\int_{\{G=0\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)$
$\leq\dfrac{\sqrt{d}\left(q+\mathcal{E}^{0}_{r}\right)^{\mu}}{\eta}+\dfrac{\sqrt{d}e^{-\alpha\left(q-\tau_{3}(r)\right)}}{\rho_{t}\left(B^{\infty}(v^{*},r)\right)}\int_{\{G=0\}}\|v-v^{*}\|_{2}\,d\rho_{t}(v).$
Here, $\mathcal{E}^{0}_{r}$, $\mathcal{E}_{\infty}$ and $\tau_{3}$ are quantities defined in Assumption 2 (C) and Lemma 4.4.
Proof.
Let $\tilde{r}=\dfrac{(q+\mathcal{E}^{0}_{r})^{\mu}}{\eta}.$ One can verify that
  1. (1)
    $\tilde{r}\geq r$
  2. (2)
    $\mathcal{E}(v)-\mathcal{E}^{0}_{r}\geq q$ $\forall v\in\{G=0\}\cap B^{\infty}(v^{*},\tilde{r})^{c}$.
For (1), we begin by computing directly:
$\tilde{r}=\dfrac{(q+\mathcal{E}^{0}_{r})^{\mu}}{\eta}\geq\dfrac{(\mathcal{E}^{0}_{r})^{\mu}}{\eta}=\dfrac{(\mathcal{E}^{0}_{r}-\underline{\mathcal{E}_{0}})^{\mu}}{\eta},$
where the last equality is because $\underline{\mathcal{E}_{0}}=\mathcal{E}(v^{*})=0.$ Then for any $v\in B^{\infty}(v^{*},r)\cap\{G=0\}$, by the definition of $\mathcal{E}^{0}_{r},$ in Lemma 4.4, one has
$\dfrac{(\mathcal{E}^{0}_{r}-\underline{\mathcal{E}_{0}})^{\mu}}{\eta}\geq\dfrac{(\mathcal{E}(v)-\underline{\mathcal{E}_{0}})^{\mu}}{\eta}.$
Then we use Assumption 2 (C2) to get
$\tilde{r}\geq\dfrac{(\mathcal{E}(v)-\underline{\mathcal{E}_{0}})^{\mu}}{\eta}\geq\big\|v-v^{*}\big\|_{\infty}.$
By Assumption 2 (C1), $\partial B^{\infty}(v^{*},r)\cap\{G=0\}\neq\emptyset$, which leads to
$\sup_{v\in B^{\infty}(v^{*},r)\cap\{G=0\}}\big\|v-v^{*}\big\|_{\infty}=r.$
Since the above inequality holds $\forall v\in B^{\infty}(v^{*},r)\cap\{G=0\}$, one then has $\tilde{r}\geq r$, which completes the proof of the first one. And for (2), for all $v\in\{G=0\}\cap B^{\infty}(v^{*},r)^{c}$, we can compute:
$\mathcal{E}(v)-\mathcal{E}^{0}_{r}$ $=\mathcal{E}(v)-\underline{\mathcal{E}_{0}}-(\mathcal{E}^{0}_{r}-\underline{\mathcal{E}_{0}})$
$\geq(\eta\big\|v-v^{*}\big\|_{\infty})^{1/\mu}-(\mathcal{E}^{0}_{r}-\underline{\mathcal{E}_{0}})\geq(\eta\tilde{r})^{1/\mu}-(\mathcal{E}^{0}_{r}-\underline{\mathcal{E}_{0}})=q+\underline{\mathcal{E}_{0}}=q,$
where the first inequality comes from Assumption 2 (C2), the second inequality is due to $v\in B^{\infty}(v^{*},\tilde{r})^{c}$, the third inequality is because of the definition of $\tilde{r}$ and the last equality is because we assumed $\mathcal{E}(v^{*})=0$. This completes the proof of the second one.
Then we have
$\int_{\{G=0\}}\dfrac{\big\|v-v^{*}\big\|_{\infty}}{\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)=$ $\int_{\{G=0\}\cap B^{\infty}(v^{*},\tilde{r})}\dfrac{\big\|v-v^{*}\big\|_{\infty}}{\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)$
$+\int_{\{G=0\}\cap B^{\infty}(v^{*},\tilde{r})^{c}}\dfrac{\big\|v-v^{*}\big\|_{\infty}}{\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v).$
For the former term, we have the following estimate
(69) $\int_{\{G=0\}\cap B^{\infty}(v^{*},\tilde{r})}\dfrac{\big\|v-v^{*}\big\|_{\infty}}{\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)$
$\leq\tilde{r}\int_{\{G=0\}\cap B^{\infty}(v^{*},\tilde{r})}\dfrac{1}{\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)\leq\tilde{r}$.
For the latter term, we first notice that
$\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}$ $=\int e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)$
$\geq\int_{B^{\infty}(v^{*},r)}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)$
$\geq\int_{B^{\infty}(v^{*},r)}e^{-\alpha\mathcal{E}_{r}}\,d\rho_{t}(v)$
$=e^{-\alpha\mathcal{E}_{r}}\rho_{t}\left(B^{\infty}(v^{*},r)\right).$
Here the second inequality is because of the definition of $\mathcal{E}_{r}$ in Lemma 4.4. So
(70) $\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}\geq e^{-\alpha\mathcal{E}_{r}}\rho_{t}\left(B^{\infty}(v^{*},r)\right)$
holds true for any choice of $\alpha$ and $r$. Then one can deduce
$\int_{\{G=0\}\cap B^{\infty}(v^{*},\tilde{r})^{c}}\dfrac{\big\|v-v^{*}\big\|_{\infty}}{\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)$
$\leq\int_{\{G=0\}\cap B^{\infty}(v^{*},\tilde{r})^{c}}\dfrac{\big\|v-v^{*}\big\|_{\infty}}{\rho_{t}\left(B^{\infty}(v^{*},r)\right)}e^{-\alpha\left(\mathcal{E}(v)-\mathcal{E}_{r}\right)}\,d\rho_{t}(v)$
$\leq\int_{\{G=0\}\cap B^{\infty}(v^{*},\tilde{r})^{c}}\dfrac{\big\|v-v^{*}\big\|_{\infty}}{\rho_{t}\left(B^{\infty}(v^{*},r)\right)}e^{-\alpha\left(\mathcal{E}(v)-\mathcal{E}^{0}_{r}-\tau_{3}(r)\right)}\,d\rho_{t}(v)$
$\leq\int_{\{G=0\}}\dfrac{\big\|v-v^{*}\big\|_{\infty}}{\rho_{t}\left(B^{\infty}(v^{*},r)\right)}e^{-\alpha\left(q-\tau_{3}(r)\right)}\,d\rho_{t}(v),$
where in the second inequality, we used Lemma 4.4 and in the third third inequality, we used the fact (2) that $\mathcal{E}(v)-\mathcal{E}^{0}_{r}\geq q$ $\forall v\in\{G=0\}\cap B^{\infty}(v^{*},\tilde{r})^{c}$. Thus
$\int_{\{G=0\}\cap B^{\infty}(v^{*},\tilde{r})^{c}}\dfrac{\big\|v-v^{*}\big\|_{\infty}}{\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)\leq\int_{\{G=0\}}\dfrac{\big\|v-v^{*}\big\|_{\infty}}{\rho_{t}\left(B^{\infty}(v^{*},r)\right)}e^{-\alpha\left(q-\tau_{3}(r)\right)}\,d\rho_{t}(v).$
Combining the above inequality and (69), we can get
$\int_{\{G=0\}}\dfrac{\big\|v-v^{*}\big\|_{\infty}}{\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)$
$\leq\dfrac{(q+\mathcal{E}^{0}_{r})^{\mu}}{\eta}+\dfrac{e^{-\alpha\left(q-\tau_{3}(r)\right)}}{\rho_{t}\left(B^{\infty}(v^{*},r)\right)}\int_{\{G=0\}}\big\|v-v^{*}\big\|_{\infty}\,d\rho_{t}(v).$
Since $\big\|\cdot\big\|_{\infty}\leq\big\|\cdot\big\|_{2}\leq\sqrt{d}\big\|\cdot\big\|_{2}$, we have
$\int_{\{G=0\}}\dfrac{\big\|v-v^{*}\big\|_{2}}{\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)$
$\leq\dfrac{\sqrt{d}(q+\mathcal{E}^{0}_{r})^{\mu}}{\eta}+\dfrac{\sqrt{d}e^{-\alpha\left(q-\tau_{3}(r)\right)}}{\rho_{t}\left(B^{\infty}(v^{*},r)\right)}\int_{\{G=0\}}\big\|v-v^{*}\big\|_{2}\,d\rho_{t}(v).$
This completes the proof. ∎
F.2. Lemma F.2 and proof
Lemma F.2.
Fix $0<u<u_{0}$ and $r>0$ small. $\forall q>0$ satisfying the condition that $q+\mathcal{E}_{r}^{\tilde{u}}-\underline{\mathcal{E}_{\tilde{u}}}$$\mathcal{E}_{\infty}$ is true $\forall\tilde{u}\in(0,u)$, then
$\int_{\{G\in(0,u)\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)$
$\leq\dfrac{\sqrt{d}\left(q+\mathcal{E}^{0}_{r}+\tau_{2}(u)+\tau_{4}(\max\{u,r\})\right)^{\mu}}{\eta}$
$\quad+\dfrac{\sqrt{d}e^{-\alpha\left(q-\tau_{3}(\max\{u,r\})\right)}}{\rho_{t}\left(B^{\infty}(v^{*},r)\right)}\int_{\{G\in(0,u)\}}\|v-v_{G(v)}\|_{2}\,d\rho_{t}(v)$
$\quad+\sqrt{d}\tau_{1}(u).$
Here, $v_{G(v)}=\arg\min_{v^{\prime}\in\{G(v^{\prime})=G(v)\}}\mathcal{E}(v^{\prime})$, $\underline{\mathcal{E}_{\tilde{u}}}$ and $\tau_{1}$ are defined in Assumption 2 (C1), $\mathcal{E}_{r}^{\tilde{u}}$, $\tau_{3}$ and $\tau_{4}$ are quantities defined in Lemma 4.4.
Proof.
We first can deduce
$\int_{\{G\in(0,u)\}}\dfrac{\big\|v-v^{*}\big\|_{\infty}}{\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)$
$=\int_{\{G\in(0,u)\}}\dfrac{\big\|v-v^{*}\big\|_{\infty}}{\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}\big\|\nabla G\big\|_{2}}e^{-\alpha\mathcal{E}(v)}\big\|\nabla G\big\|_{2}\rho_{t}\,dv$
$=\int_{0}^{u}\,d\tilde{u}\int_{\{G(v)=\tilde{u}\}}\dfrac{\big\|v-v^{*}\big\|_{\infty}}{\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}\big\|\nabla G\big\|_{2}}e^{-\alpha\mathcal{E}(v)}\rho_{t}\,dH_{d-1}(v).$
Here, the first equality is because of Assumption 2 (B3) that $\nabla G\neq 0$ and the second equality comes from the co-area formula. $dH_{d-1}(v)$ is the (d1)(d-1) dimensional Hausdorff measure.
Now we fix $0<\tilde{u}<u$ and study the inner integral. We pick $\tilde{r}_{\tilde{u}}=\dfrac{(q+\mathcal{E}^{\tilde{u}}_{r}-\underline{\mathcal{E}_{\tilde{u}}})^{\mu}}{\eta}$. One can easily use Assumption 2 (C2) to verify the following facts:
  1. (1)
    $\tilde{r}_{\tilde{u}}\geq r$.
  2. (2)
    $\mathcal{E}(v)-\mathcal{E}^{\tilde{u}}_{r}\geq q$ for $v\in B^{\infty}(v_{\tilde{u}},\tilde{r}_{\tilde{u}})^{c}\cap\{G(v)=\tilde{u}\}$.
  3. (3)
    $\tilde{r}_{\tilde{u}}\leq\tilde{r}=\dfrac{\left(q+\mathcal{E}^{0}_{r}+\tau_{2}(u)+\tau_{4}(\max\{u,r\})\right)^{\mu}}{\eta}.$
For the proof of the first two facts, one can use the same method we used at the beginning of the proof of Lemma F.1 and details are omitted. For (3), one can prove it as follows:
$\tilde{r}_{\tilde{u}}$ $=\dfrac{\left(q+\mathcal{E}^{\tilde{u}}_{r}-\underline{\mathcal{E}_{\tilde{u}}}\right)^{\mu}}{\eta}$
$=\dfrac{\left(q+\mathcal{E}^{0}_{r}+(\mathcal{E}^{\tilde{u}}_{r}-\mathcal{E}^{0}_{r})-\underline{\mathcal{E}_{\tilde{u}}}\right)^{\mu}}{\eta}$
$\leq\dfrac{\left(q+\mathcal{E}^{0}_{r}+\tau_{4}(\max\{\tilde{u},r\})-\tau_{2}(\tilde{u})\right)^{\mu}}{\eta}\leq\dfrac{\left(q+\mathcal{E}^{0}_{r}+\tau_{4}(\max\{u,r\})-\tau_{2}(u)\right)^{\mu}}{\eta},$
where the two inequalities are because of Assumption 2 (C2) and Lemma 4.4.
Then by the triangle inequality, one obtains
$\int_{\{G(v)=\tilde{u}\}}\dfrac{\big\|v-v^{*}\big\|_{\infty}}{\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}\big\|\nabla G\big\|_{2}}e^{-\alpha\mathcal{E}(v)}\rho_{t}\,dH_{d-1}(v)$
$\leq$ $\int_{\{G(v)=\tilde{u}\}\cap B^{\infty}(v_{\tilde{u}},\tilde{r}_{\tilde{u}})}\dfrac{\big\|v-v_{\tilde{u}}\big\|_{\infty}}{\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}\big\|\nabla G\big\|_{2}}e^{-\alpha\mathcal{E}(v)}\rho_{t}\,dH_{d-1}(v)$
$+\int_{\{G(v)=\tilde{u}\}\cap B^{\infty}(v_{\tilde{u}},\tilde{r}_{\tilde{u}})^{c}}\dfrac{\big\|v-v_{\tilde{u}}\big\|_{\infty}}{\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}\big\|\nabla G\big\|_{2}}e^{-\alpha\mathcal{E}(v)}\rho_{t}\,dH_{d-1}(v)$
$+\int_{\{G(v)=\tilde{u}\}}\dfrac{\big\|v^{*}-v_{\tilde{u}}\big\|_{\infty}}{\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}\big\|\nabla G\big\|_{2}}e^{-\alpha\mathcal{E}(v)}\rho_{t}\,dH_{d-1}(v).$
Thus one needs to bound the above three terms. For the first one,
$\int_{\{G(v)=\tilde{u}\}\cap B^{\infty}(v_{\tilde{u}},\tilde{r}_{\tilde{u}})}\dfrac{\big\|v-v_{\tilde{u}}\big\|_{\infty}}{\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}\big\|\nabla G\big\|_{2}}e^{-\alpha\mathcal{E}(v)}\rho_{t}\,dH_{d-1}(v)$
$\leq\tilde{r}_{\tilde{u}}\int_{\{G(v)=\tilde{u}\}}\dfrac{1}{\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}\big\|\nabla G\big\|_{2}}e^{-\alpha\mathcal{E}(v)}\rho_{t}\,dH_{d-1}(v)$
$\leq\tilde{r}\int_{\{G(v)=\tilde{u}\}}\dfrac{1}{\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}\big\|\nabla G\big\|_{2}}e^{-\alpha\mathcal{E}(v)}\rho_{t}\,dH_{d-1}(v).$
For the second one,
$\int_{\{G(v)=\tilde{u}\}\cap B^{\infty}(v_{\tilde{u}},\tilde{r}_{\tilde{u}})^{c}}\dfrac{\big\|v-v_{\tilde{u}}\big\|_{\infty}}{\big\|\omega_{\alpha}\big\|_{1}\big\|\nabla G\big\|_{2}}e^{-\alpha\mathcal{E}(v)}\rho_{t}\,dH_{d-1}(v)$
$\leq\int_{\{G(v)=\tilde{u}\}\cap B^{\infty}(v_{\tilde{u}},\tilde{r}_{\tilde{u}})^{c}}\dfrac{\big\|v-v_{\tilde{u}}\big\|_{\infty}}{\rho_{t}\left(B^{\infty}(v^{*},r)\right)\big\|\nabla G\big\|_{2}}e^{-\alpha\left(\mathcal{E}(v)-\mathcal{E}_{r}\right)}\rho_{t}\,dH_{d-1}(v)$
$\leq\int_{\{G(v)=\tilde{u}\}\cap B^{\infty}(v_{\tilde{u}},\tilde{r}_{\tilde{u}})^{c}}\dfrac{\big\|v-v_{\tilde{u}}\big\|_{\infty}}{\rho_{t}\left(B^{\infty}(v^{*},r)\right)\big\|\nabla G\big\|_{2}}e^{-\alpha\left(\mathcal{E}(v)-\mathcal{E}^{\tilde{u}}_{r}-\tau_{3}(\max\{\tilde{u},r\})\right)}\rho_{t}\,dH_{d-1}(v)$
$\leq\int_{\{G(v)=\tilde{u}\}\cap B^{\infty}(v_{\tilde{u}},\tilde{r}_{\tilde{u}})^{c}}\dfrac{\big\|v-v_{\tilde{u}}\big\|_{\infty}}{\rho_{t}\left(B^{\infty}(v^{*},r)\right)\big\|\nabla G\big\|_{2}}e^{-\alpha\left(\mathcal{E}(v)-\mathcal{E}^{\tilde{u}}_{r}-\tau_{3}(\max\{u,r\})\right)}\rho_{t}\,dH_{d-1}(v)$
$\leq\int_{\{G(v)=\tilde{u}\}\cap B^{\infty}(v_{\tilde{u}},\tilde{r}_{\tilde{u}})^{c}}\dfrac{\big\|v-v_{\tilde{u}}\big\|_{\infty}}{\rho_{t}\left(B^{\infty}(v^{*},r)\right)\big\|\nabla G\big\|_{2}}e^{-\alpha\left(q-\tau_{3}(\max\{u,r\})\right)}\rho_{t}\,dH_{d-1}(v)$
$\leq\dfrac{e^{-\alpha\left(q-\tau_{3}(\max\{u,r\})\right)}}{\rho_{t}\left(B^{\infty}(v^{*},r)\right)}\int_{\{G(v)=\tilde{u}\}}\dfrac{\big\|v-v_{\tilde{u}}\big\|_{\infty}}{\big\|\nabla G\big\|_{2}}\rho_{t}\,dH_{d-1}(v),$
where in the first inequality above, we used (70) and in the second and third inequalities above, we used Lemma 4.4 that $|\mathcal{E}_{r}^{u}-\mathcal{E}_{r}|\leq\tau_{3}(\max\{u,r\})$ and the assumption that $\tau_{3}$ is an increasing function. In the fourth inequality, we used the fact (2) that $\mathcal{E}(v)-\mathcal{E}^{\tilde{u}}_{r}\geq q$ for $v\in B(v_{\tilde{u}},\tilde{r}_{\tilde{u}})^{c}\cap\{G(v)=\tilde{u}\}$.
For the third term,
$\int_{\{G(v)=\tilde{u}\}}\dfrac{\big\|v^{*}-v_{\tilde{u}}\big\|_{\infty}}{\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}\big\|\nabla G\big\|_{2}}e^{-\alpha\mathcal{E}(v)}\rho_{t}\,dH_{d-1}(v)$
$=\big\|v^{*}-v_{\tilde{u}}\big\|_{\infty}\int_{\{G(v)=\tilde{u}\}}\dfrac{1}{\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}\big\|\nabla G\big\|_{2}}e^{-\alpha\mathcal{E}(v)}\rho_{t}\,dH_{d-1}(v)$
$\leq\tau_{1}(\tilde{u})\int_{\{G(v)=\tilde{u}\}}\dfrac{1}{\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}\big\|\nabla G\big\|_{2}}e^{-\alpha\mathcal{E}(v)}\rho_{t}\,dH_{d-1}(v)$
$\leq\tau_{1}(u)\int_{\{G(v)=\tilde{u}\}}\dfrac{1}{\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}\big\|\nabla G\big\|_{2}}e^{-\alpha\mathcal{E}(v)}\rho_{t}\,dH_{d-1}(v),$
where in the first and second inequalities, we used Assumption 2 (C1) that $\big\|v_{u}-v^{*}\big\|\leq\tau_{1}(u)$ and the fact that $\tau_{1}$ is an increasing function. Thus
$\int_{\{G(v)=\tilde{u}\}}\dfrac{\big\|v-v^{*}\big\|_{\infty}}{\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}\big\|\nabla G\big\|_{2}}e^{-\alpha\mathcal{E}(v)}\rho_{t}\,dH_{d-1}(v)$
$\leq$ $\tilde{r}\int_{\{G(v)=\tilde{u}\}}\dfrac{1}{\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}\big\|\nabla G\big\|_{2}}e^{-\alpha\mathcal{E}(v)}\rho_{t}\,dH_{d-1}(v)$
$+\dfrac{e^{-\alpha\left(q-\tau_{3}(\max\{u,r\})\right)}}{\rho_{t}\left(B^{\infty}(v^{*},r)\right)}\int_{\{G(v)=\tilde{u}\}}\dfrac{\big\|v-v_{\tilde{u}}\big\|_{\infty}}{\big\|\nabla G\big\|_{2}}\rho_{t}\,dH_{d-1}(v)$
$+\tau_{1}(u)\int_{\{G(v)=\tilde{u}\}}\dfrac{1}{\big\|\omega_{\alpha}\big\|_{1}\big\|\nabla G\big\|_{2}}e^{-\alpha\mathcal{E}(v)}\rho_{t}\,dH_{d-1}(v).$
We can integrate the above inequality with respect to $\tilde{u}$ from 0 to $u$ to get
$\int_{\{G\in(0,u)\}}\dfrac{\big\|v-v^{*}\big\|_{\infty}}{\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)$
$\leq$ $\left(\tilde{r}+\tau_{1}(u)\right)\int_{0}^{u}\,d\tilde{u}\int_{\{G(v)=\tilde{u}\}}\dfrac{1}{\big\|\omega_{\alpha}\big\|_{L^{1}(\rho_{t})}\big\|\nabla G\big\|_{2}}e^{-\alpha\mathcal{E}(v)}\rho_{t}\,dH_{d-1}(v)$
$+\dfrac{e^{-\alpha\left(q-\tau_{3}(\max\{u,r\})\right)}}{\rho_{t}\left(B^{\infty}(v^{*},r)\right)}\int_{0}^{u}\,d\tilde{u}\int_{\{G(v)=\tilde{u}\}}\dfrac{\big\|v-v_{\tilde{u}}\big\|_{\infty}}{\big\|\nabla G\big\|_{2}}\rho_{t}\,dH_{d-1}(v)$
$=$ $\tilde{r}+\tau_{1}(u)+\dfrac{e^{-\alpha\left(q-\tau_{3}(\max\{u,r\})\right)}}{\rho_{t}\left(B^{\infty}(v^{*},r)\right)}\int_{\{G\in(0,u)\}}\big\|v-v_{G(v)}\big\|_{\infty}\,d\rho_{t}(v),$
where in the equality, we used the co-area formula again and the definition of $\omega_{\alpha}$. Then combining with the fact that $\big\|\cdot\big\|_{\infty}\leq\big\|\cdot\big\|_{2}\leq\sqrt{d}\big\|\cdot\big\|_{2}$ finishes the proof. ∎
Appendix G Proof of Laplace Principle: Lemma 4.5
In this section, we present the proof of Lemma 4.5.
Proof.
By the definition of the consensus point $v_{\alpha}\big(\rho_{t}\big)$, one has
$\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{2}$ $=\|\int v\cdot\dfrac{e^{-\alpha\mathcal{E}(v)}}{\|\omega_{\alpha}\|_{1}}\,d\rho_{t}(v)-v^{*}\|_{2}$
$=\|\int(v-v^{*})\cdot\dfrac{e^{-\alpha\mathcal{E}(v)}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{t})}}\,d\rho_{t}(v)\|_{2}\leq\int\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v),$
where we used Minkowski’s inequality. Then we can compute:
$\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\|_{2}$
$\leq\int\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)$
$=\int_{\{G=0\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)$
$\quad+\int_{\{G\in(0,u)\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)$
$\quad+\int_{\{G\geq u\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v).$
For the first term, we can upper bound it using Lemma F.1:
$\int_{\{G=0\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)$
$\leq\dfrac{\sqrt{d}(q+\mathcal{E}^{0}_{r})^{\mu}}{\eta}+\dfrac{\sqrt{d}e^{-\alpha\left(q-\tau_{3}(r)\right)}}{\rho_{t}\left(B^{\infty}(v^{*},r)\right)}\int_{\{G=0\}}\|v-v^{*}\|_{2}\,d\rho_{t}(v)$
$\leq\dfrac{\sqrt{d}\left(q+\mathcal{E}^{0}_{r}+\tau_{2}(u)+\tau_{4}(\max\{u,r\})\right)^{\mu}}{\eta}$
$\quad+\dfrac{\sqrt{d}e^{-\alpha\left(q-\tau_{3}(r)\right)}}{\rho_{t}\left(B^{\infty}(v^{*},r)\right)}\int_{\{G=0\}}\|v-v^{*}\|_{2}\,d\rho_{t}(v).$
For the second term, we can upper bound it using Lemma F.2:
$\int_{\{G\in(0,u)\}}\dfrac{\|v-v^{*}\|_{2}}{\|\omega_{\alpha}\|_{L^{1}(\rho_{t})}}e^{-\alpha\mathcal{E}(v)}\,d\rho_{t}(v)$
$\leq\dfrac{\sqrt{d}\left(q+\mathcal{E}^{0}_{r}+\tau_{2}(u)+\tau_{4}(\max\{u,r\})\right)^{\mu}}{\eta}$
$\quad+\dfrac{\sqrt{d}e^{-\alpha\left(q-\tau_{3}(\max\{u,r\})\right)}}{\rho_{t}\left(B^{\infty}(v^{*},r)\right)}\int_{\{G\in(0,u)\}}\|v-v_{G(v)}\|_{2}\,d\rho_{t}(v)$
$\quad+\sqrt{d}\tau_{1}(u).$
Finally, We leave the third term unchanged. Combining the estimates for the above three terms, we can finish the proof. ∎
Appendix H The Complete Proof of Lemma 4.6
Proof.
Since $\phi_{r}\leq 1$, one can show that
$\rho_{t}\left(B(v^{*},r)\right)\geq\int\phi_{r}(v)\,d\rho_{t}(v).$
So it suffices to find a lower bound for $\int\phi_{r}(v)\,d\rho_{t}(v).$ To do this, since $\phi_{r}\in\mathcal{C}_{*}^{2}(\mathbb{R}^{d})$, one can plug $\phi_{r}$ into the Definition (3.2) to get that
$\dfrac{d}{dt}\int\phi_{r}(v)\,d\rho_{t}(v)=\int\left(T_{1}(v)+T_{2}(v)+T_{3}(v)\right)\,d\rho_{t}(v),$
where
$T_{1}(v)=-\lambda\left(v-v_{\alpha}\big(\rho_{t}\big)\right)\cdot\nabla\phi_{r}(v),$
$T_{2}(v)=\dfrac{\sigma^{2}}{2}\sum_{k=1}^{d}\left(v-v_{\alpha}\big(\rho_{t}\big)\right)_{k}^{2}\partial_{kk}\phi_{r}(v)$
and
$T_{3}(v)=-\dfrac{1}{\epsilon}\int\Bigl\langle\nabla G,\nabla\phi\Bigr\rangle.$
One can calculate directly that
$\nabla\phi_{r}(v)=-2r^{2}\dfrac{v-v^{*}}{\left(r^{2}-\big\|v-v^{*}\big\|^{2}\right)^{2}}\phi_{r}(v),$
$\partial_{kk}\phi_{r}(v)=2r^{2}\left(\dfrac{2\left(2(v-v^{*})_{k}^{2}-r^{2}\right)(v-v^{*})_{k}^{2}-d\left(r^{2}-(v-v^{*})_{k}^{2}\right)^{2}}{\left(r^{2}-(v-v^{*})_{k}^{2}\right)^{4}}\right)\phi_{r}(v).$
By the expression of $\nabla\phi_{r}$, one knows that $T_{3}\geq 0$ because of Assumption 2 (B1). Thus wone only has to find the lower bound of $T_{1}$ and $T_{2}$. The details of bounding them are exactly the same as Proposition 2 [22]. Following the same steps, it turns out
$\int\left(T_{1}(v)+T_{2}(v)\right)\,d\rho_{t}(v)\geq-a\int\phi_{r}(v)\,d\rho_{t}(v),$
where $a$ is the constant defined in the statement of Theorem 4.6. Thus
$\dfrac{d}{dt}\int\phi_{r}(v)\,d\rho_{t}(v)$
$=\int\left(T_{1}(v)+T_{2}(v)+T_{3}(v)\right)\,d\rho_{t}(v)$
$\geq\int\left(T_{1}(v)+T_{2}(v)\right)\,d\rho_{t}(v)\geq-a\int\phi_{r}(v)\,d\rho_{t}(v).$
Then applying Gronwall’s inequality will finish the proof. ∎
Appendix I Proof of Lemma 4.7
Proof.
Let $B=\sup_{t\in[0,T]}\big\|v_{\alpha}\big(\rho_{t}\big)-v^{*}\big\|_{2}$ and $\tilde{B}=\sup_{t\in[0,T]}\mathcal{V}\big(\rho_{t}\big)$. Also, because of Assumption 2 (B2) that $G(v)\in C^{2}_{*}(\mathbb{R}^{d})$ and $G(v)\lesssim\big\|\nabla G(v)\big\|_{2}^{2}$, one can find some positive constant $\tilde{c}$ such that
(71) $|\partial_{kk}G(v)|\leq\tilde{c}\text{, }\big\|\nabla G(v)\big\|\leq\tilde{c}\left(1+\big\|v-v^{*}\big\|\right)$
and
(72) $G(v)\leq\tilde{c}\big\|\nabla G(v)\big\|^{2}.$
Plug $G$ into Definition 3.2 gives
$\dfrac{d}{dt}\int G\,d\rho_{t}(v)$ $=$ $-\lambda\int\Bigl\langle v-v_{\alpha}\big(\rho_{t}\big),\nabla G\Bigr\rangle\,d\rho_{t}(v)$
$+\dfrac{\sigma^{2}}{2}\int\sum_{k=1}^{d}\left(v-v_{\alpha}\big(\rho_{t}\big)\right)_{k}^{2}\partial_{kk}G\,d\rho_{t}(v)-\dfrac{1}{\epsilon}\int\big\|\nabla G\big\|_{2}^{2}\,d\rho_{t}(v)$
$\leq$ $-\lambda\int\Bigl\langle v-v^{*},\nabla G\Bigr\rangle\,d\rho_{t}(v)-\lambda\int\Bigl\langle v^{*}-v_{\alpha}\big(\rho_{t}\big),\nabla G\Bigr\rangle\,d\rho_{t}(v)$
$+\sigma^{2}\int\sum_{k=1}^{d}\left((v-v^{*})_{k}^{2}+\left(v^{*}-v_{\alpha}\big(\rho_{t}\big)\right)_{k}^{2}\right)\partial_{kk}G\,d\rho_{t}(v)$
$-\dfrac{1}{\epsilon}\int\big\|\nabla G\big\|_{2}^{2}\,d\rho_{t}(v).$
The first term is non-positive because of Assumption 2 (B1) and the second term can be bounded as follows
$-\lambda\int\Bigl\langle v^{*}-v_{\alpha}\big(\rho_{t}\big),\nabla G\Bigr\rangle\,d\rho_{t}(v)$ $\leq\lambda\int\big\|v^{*}-v_{\alpha}\big(\rho_{t}\big)\big\|\big\|\nabla G\big\|\,d\rho_{t}(v)$
$\leq\lambda\int B\big\|\nabla G\big\|\,d\rho_{t}(v)$
$\leq\lambda B\tilde{c}\int\left(1+\big\|v-v^{*}\big\|\right)\,d\rho_{t}(v)$
[…]
where the third inequality above is due to (71). The third term is bounded above by $\tilde{c}\sigma^{2}(\tilde{B}+B^{2})$ and the fourth term is upper bounded by $\dfrac{1}{\tilde{c}\epsilon}\int G\,d\rho_{t}(v)$ because of (72). Thus one has
$\dfrac{d}{dt}\int G\,d\rho_{t}(v)\leq\lambda B\tilde{c}\left(1+\sqrt{2\tilde{B}}\right)+\tilde{c}\sigma^{2}\left(\tilde{B}+B^{2}\right)-\dfrac{1}{\epsilon\tilde{c}}\int G\,d\rho_{t}(v).$
We use $D$ to denote $\lambda B\tilde{c}\left(1+\sqrt{2\tilde{B}}\right)+\tilde{c}\sigma^{2}(\tilde{B}+B^{2}).$
Now consider $f$ satisfying
$\dfrac{d}{dt}f=D-\dfrac{1}{\tilde{c}\epsilon}f$
with initial condition $f(0)=\int G\,d\rho_{0}(v)$. By the comparison theorem, one knows that before $T$, $\int G\,d\rho_{t}(v)$ is dominated by $f$, i.e., $\int G\,d\rho_{t}(v)\leq f(t)$. And one has an explicit expression for $f$:
$f(t)=\tilde{c}\epsilon D+(\int G\,d\rho_{0}(v)-\tilde{c}\epsilon D)e^{-(1/\tilde{c}\epsilon)t}.$
When $\epsilon$ is small enough, i.e.,
(73) $\epsilon<\frac{\int G\,d\rho_{0}(v)}{\tilde{c}D},$
one can deduce
$f(t)$ $=\tilde{c}\epsilon D+\left(\int G\,d\rho_{0}(v)-\tilde{c}\epsilon D\right)e^{-(1/\tilde{c}\epsilon)t}$
$\leq\tilde{c}\epsilon D+\left(\int G\,d\rho_{0}(v)-\tilde{c}\epsilon D\right)=\int G\,d\rho_{0}(v).$
Thus for $t\in[0,T]$,
$\int G\,d\rho_{t}(v)\leq\int G\,d\rho_{0}(v).$
This completes the proof. ∎
Appendix J Computational Details in the Proof of Theorem 4.1: Verification of the Assumptions in Lemma 4.5
Proof.
To see this, firstly, by the choice of $q,r,u$, for any $\tilde{u}\in[0,u)$:
$|\underline{\mathcal{E}_{\tilde{u}}}|\leq\tilde{u}<u\leq\dfrac{1}{4}q$
and
$\mathcal{E}_{r}^{\tilde{u}}=\mathcal{E}^{0}_{r}+(\mathcal{E}_{r}^{\tilde{u}}-\mathcal{E}^{0}_{r})\leq\dfrac{1}{4}q+\max\{\tilde{u},r\}\leq\dfrac{1}{4}q+\max\{u,r\}\leq\dfrac{1}{4}q+\dfrac{1}{4}q=\dfrac{1}{2}q.$
Here the first inequality is due to the definition of $r$ and Lemma 4.4. Thus, one has
$q+\mathcal{E}_{r}^{\tilde{u}}-\underline{\mathcal{E}_{\tilde{u}}}\leq q+\dfrac{1}{2}q+\dfrac{1}{4}q=\dfrac{7}{4}q\leq\dfrac{7}{8}\mathcal{E}_{\infty}<\mathcal{E}_{\infty}.$
This verifies the assumptions in Lemma 4.5. ∎
Appendix K Details of the Numerical Experiments
K.1. Figures 1 and Figure 2
The objective function $\mathcal{E}(v)$ is the similar to (36)
$\min_{v}\quad-A\exp\left(-a\sqrt{\frac{b^{2}}{d}\left\lVert v-\hat{v}\right\rVert_{2}^{2}}\right)-\exp\left(\frac{1}{d}\sum_{i=1}^{d}\cos(2\pi b(v-\hat{v})_{i})\right)+e^{1}+A;$
with $b=3,A=20,a=0.2$. The circular constraint reads,
$g_{1}(v)=\left\lVert v\right\rVert_{2}^{2}-1;$
and the parabolic constraint reads,
$g_{2}(v)=v_{1}^{2}-v_{2}.$
The first case is a circular constraint, and the unconstrained minimizer is the same as the constrained minimizer.
$\hat{v}=v^{*}=\frac{1}{\sqrt{2}}(1,-1).$
the second case is a circular constraint, and the unconstrained minimizer is different from the constrained minimizer. Therefore,
$\hat{v}=(1/2,1/3),\quad v^{*}=(0.781475;\sqrt{1-0.781475^{2}}).$
The third case is a parabolic constraint, and the unconstrained minimizer is different from the constrained minimizer. Therefore,
$\hat{v}=(1/2,1/3),\quad v^{*}=(0.5428;0.5428^{2}).$
We use Algorithm 1 with
(74) $N=50,\ \alpha=30,\ \epsilon=0.01,\ \lambda=1,\ \sigma=1,\ \gamma=0.01,\ \epsilon_{\text{stop}}=0.$
[…]
We use Algorithm 1 in [20] for the projected CBO method. For the penalized CBO method, we set the penalty as $\frac{1}{\epsilon}G(v)$, and then apply the CBO algorithm to the following unconstrained optimization problem,
$\mathcal{E}^{\epsilon}(v)=\mathcal{E}(v)+\frac{1}{\epsilon}G(v).$
We use the same parameters as (74) for the two alternative algorithms.
Acknowledgements
JAC was supported by the Advanced Grant Nonlocal-CPD (Nonlocal PDEs for Complex Particle Dynamics: Phase Transitions, Patterns and Synchronization) of the European Research Council Executive Agency (ERC) under the European Union’s Horizon 2020 research and innovation programme (grant agreement No. 883363). JAC was also partially supported by the Engineering and Physical Sciences Research Council (EPSRC) under grants EP/T022132/1 and EP/V051121/1.
SJ was partially supported by funding XDA25010404, the NSFC grants Nos. 12350710181 and 92270001, the Shanghai Municipal Science and Technology Major Projects (Nos. 2021SHZDZX0102 and 23JC1402300), the Shanghai Municipal Science and Technology Key Project (No. 22JC1401500), and the Fundamental Research Funds for the Central Universities.
YZ was supported by the NSF grants No 2529107.
This work was started while JAC and YZ were attending a workshop at the American Institute of Mathematics in 2018 and then developed further while JAC and YZ were visiting the Simons Institute to participate in the program “Geometric Methods in Optimization and Sampling” during the Fall of 2021.